### v1

# Constraining Participation: Affordances of Feedback Features in Interfaces to Large Language Models

Abstract.
Large language models (LLMs) are now accessible to anyone with a computer, a web browser, and an internet connection via browser-based interfaces, shifting the dynamics of participation in AI development. This paper examines the affordances of interactive feedback features in ChatGPT’s interface, analysing how they shape user input and participation in LLM iteration. Drawing on a survey of ChatGPT users and applying the mechanisms and conditions framework of affordances, we demonstrate that these features encourage simple, frequent, and performance-focused feedback while discouraging collective input and discussions among users. We argue that this feedback format significantly constrains user participation, reinforcing power imbalances between users, the public, and companies developing LLMs. Our analysis contributes to the growing body of literature on participatory AI by critically examining the limitations of existing feedback processes and proposing directions for their redesign. To enable more meaningful public participation in AI development, we advocate for a shift away from processes focused on aligning model outputs with specific user preferences. Instead, we emphasise the need for processes that facilitate dialogue between companies and diverse ‘publics’ about the purpose and applications of LLMs. This approach requires attention to the ongoing work of infrastructuring—creating and sustaining the social, technical, and institutional structures necessary to address matters of concern to groups impacted by AI development and deployment.
Large Language Models, Human-AI Interaction, User Feedback, Participatory AI
copyright: none
1. Introduction
The deployment of ChatGPT by OpenAI in 2022 marked a significant shift in how large language models (LLMs) are made available to the public. ChatGPT’s browser-based interface allows broad access to interactions with LLMs through text prompts, expanding user participation in the iterative development of LLMs. Previously, companies such as OpenAI and Google took a more cautious approach. OpenAI used application programming interfaces (APIs) and an iterative approach to deploy the GPT-2 and GPT-3 family of models (Solaiman et al., 2019; Brundage et al., 2022), limiting access to developers and researchers. Similarly, Google announced versions of the Language Model for Dialogue Applications (LaMDA) in 2021 and 2022 but did not make it publicly available through an interface or model release (Thoppilan et al., 2022). Following ChatGPT’s deployment, numerous other LLMs were made accessible via browser-based interfaces (e.g., Bard/Gemini and Claude.ai) or open releases (e.g., LLaMA).
This shift in the approach to deployment has precipitated a change in the relationship between companies developing LLMs and their users—we are not only users of LLMs but also participants in their iteration. For many years, users and non-users have been represented in datasets used to train models, constituting a form of passive and involuntary participation (Sloane et al., 2022). However, with recent deployments of public-facing interfaces, users now have a more active and voluntary avenue for participation by submitting feedback on prompt completions through interactive features in these interfaces.
This article focuses on the affordances of interactive feedback features in public-facing interfaces to LLMs, using ChatGPT as a case study. Feedback received through these features plays a role in the iteration of LLMs by aligning model outputs with specific preferences. This “bottom-up” approach to alignment elicits preferences from broad groups of users (Gabriel, 2020). Others have sought to expand the range of perspectives represented in these groups beyond the limited set of human raters involved in instruction-tuning LLMs before deployment (Kirk et al., 2024). However, our analysis focuses on how the affordances of the post-deployment feedback system amplify certain perspectives and concerns while marginalising others. Specifically, we address two research questions:
  1. (1)
    What are the affordances of interactive feedback features in the ChatGPT interface?
  2. (2)
    How do those affordances shape the nature and scope of user input and participation in the iteration of LLMs?
To address these questions, we first surveyed ChatGPT users who have submitted feedback. We then drew on the survey insights to analyse interactive features in the ChatGPT interface that allow users to submit feedback, using Davis’s (Davis, 2020) mechanisms and conditions framework of affordances. Our analysis indicates that the ChatGPT interface encourages simple, frequent, individualised, and unidirectional feedback on performance-related concerns while discouraging feedback from collectives or discussions among users.
We discuss implications for (re)designing feedback processes for LLMs. First, we explore deliberative approaches that might enable bidirectional communication, changes in perspectives over time, and discussion about social or ethical issues among groups of users. However, we recognise that deliberation on rulesets for model outputs will not fundamentally alter the power dynamics between users, the public, and platforms. Second, we consider the “infrastructuring” (Le Dantec and DiSalvo, 2013) required to support and sustain more expansive feedback channels between the public and technology companies. We conclude by urging companies deploying LLMs to shift their focus from aligning model outputs with user preferences to fostering sustained dialogue with diverse ‘publics’ about the purposes and applications of these technologies. This approach requires careful attention to the ongoing work of infrastructuring participation—creating and maintaining the social, technical, and institutional structures necessary for genuine public engagement in AI development. We must move beyond preference elicitation before LLMs become ubiquitous infrastructure—systems that are embedded in other structures, transparent in use, and only visible upon breakdown (Star and Ruhleder, 1996).
2. Background
In this section, we explore concepts related to participatory approaches, feedback features, and the challenges of scale in the context of LLMs. First, we provide an overview of the alignment problem and the “participatory turn” in AI (Delgado et al., 2023). Next, we examine feedback features used in recent deployments of LLMs through browser-based interfaces. Then, we explore the implications of scale for the “sameness” of those features across diverse contexts (Young et al., 2024). We conclude the section by proposing the study of these concepts through the mechanisms and conditions framework of affordances.
2.1. Alignment and Participatory Approaches
In recent years, technology companies have explored various forms of user participation to address the alignment problem—ensuring that AI systems behave in ways that align with human values (Russell, 2019). Two fundamental questions pertinent to the alignment problem are: what values should AI systems align with, and who should decide those values? (Gabriel, 2020) One approach to addressing both questions is bottom-up—eliciting preferences from a broadly defined public and inferring from those preferences the values to align with (Gabriel, 2020).
Interest in participatory approaches for aligning LLMs is part of a broader participatory turn in AI research, design, and development (Delgado et al., 2023). Participatory Al projects involve end users and those impacted by AI systems in the design and development of those systems (Cooper and Zafiroglu, 2024), drawing on long-standing traditions in other fields, such as participatory design (Muller and Kuhn, 1993; Sundblad, 2011) and participatory planning (Arnstein, 1969; Kelty, 2020). The turn to participation in AI research, design and development has included papers, panels and workshops on participatory machine learning (ML) at prominent computer science conferences (Kulynych et al., 2020; Zytko et al., 2022) and calls for public participation in AI in recent policy frameworks (Farthing et al., 2021; U.S. Office of Science and Technology Policy, 2022; Tabassi, 2023).
Participatory approaches promise more “democratic” AI systems (Kulynych et al., 2020) and more “pluralistic” alignment (Kirk et al., 2024). However, there is a broad spectrum of participatory practices in AI design, from informing or consulting participants to equitable partnerships with participants (Birhane et al., 2022; Corbett et al., 2023; Delgado et al., 2023). This spectrum of participatory practices also existed in participatory design, participatory planning, and other antecedent fields. For the development of large-scale AI systems, participation has often taken passive and involuntary forms through the representation of groups of individuals in datasets used to train models (Denton et al., 2020; Sloane et al., 2022). While feedback features in interfaces to LLMs provide a more active and voluntary form of user participation, their emergence as the primary mode for the public to influence development processes has raised questions about the conceptual leap from structured preference elicitation to genuine participatory engagement (Feffer et al., 2023b).
Additionally, there is a significant gap between the scale of participatory methods and the training and deployment of recent AI systems (Diaz and Madaio, 2023). Surveys of participatory approaches in computing research (Cooper et al., 2022) and Participatory AI projects (Delgado et al., 2023) indicate that engagements have primarily been localised, focusing on one or a few groups or communities with shared characteristics. In contrast, recent deployments of LLMs through browser-based interfaces are accessible to anyone with access to a computer, a web browser, and an internet connection.
2.2. Interactive Feedback Features in Interfaces to LLMs
Feedback has a rich history rooted in cybernetics, which introduced the idea of feedback loops, where the output of a system is fed back as input, and queried the implications of feedback loops for modelling dynamical systems over time (Wiener, 1948). The concept of feedback loops influenced subsequent AI research directions (Minsky, 1961) and remains fundamental to AI systems that learn over time. Concurrently, ‘feedback’ has also become a general term referring to information provided by an observer (e.g., users or other stakeholders) assessing a system’s performance.
Reinforcement Learning from Human Feedback (RLHF) demonstrates how the two types of feedback are combined in modern AI development. RLHF begins with evaluative feedback, when a model generates two responses to a given prompt and human raters determine which response is better than the other according to criteria. These human assessments are then used to create a cybernetic feedback loop: an RL algorithm uses the assessments to predict how a human would rate other responses, and an LLM is fine-tuned on the resulting reward model (Ziegler et al., 2019; Ouyang et al., 2022; Bai et al., 2022). This approach has become central to aligning pre-trained LLMs with specific preferences (Lambert et al., 2023). However, integrating human feedback into the learning process increases the time to train a model and the labour costs involved (Maslej et al., 2024). RLHF requires large, diverse groups of human raters to assess prompt completions before fine-tuning the model in the final stage (Prabhakaran et al., 2021).
The importance of RLHF for aligning LLMs has increased the demand for human feedback data (Lambert et al., 2023) and research attention to practices for creating and curating diverse, representative, and quality human feedback data (Feffer et al., 2023a; Kirk et al., 2024; Prabhakaran et al., 2021). However, less attention has been paid to human feedback collected through interactive features in interfaces to LLMs, in which users evaluate prompt completions of deployed systems. This form of feedback is captured after a given LLM has been trained and deployed, and such feedback may be used to improve future model iterations. For example, OpenAI used feedback on non-factual responses obtained through the feedback features in ChatGPT’s interface, with additional labelled comparison data, to train reward models that reduced GPT-4’s tendency to “hallucinate” compared to GPT-3.5 (OpenAI, 2023).
There is growing interest in crowdsourcing human feedback datasets to align LLMs with user preferences (e.g., LeCun, 2023). In the blog post announcing the deployment of ChatGPT, OpenAI emphasised their interest in collecting user feedback to uncover potential risks and harms:
“We are … interested in feedback regarding harmful outputs that could occur in real-world, non-adversarial conditions, as well as feedback that helps us uncover and understand novel risks and possible mitigations.” (OpenAI, 2022)
OpenAI launched a feedback competition alongside the ChatGPT interface to encourage user participation (running through December 2022). Users were encouraged to submit free-form feedback via a Google form that could help OpenAI identify potential real-world risks and harms, with winners receiving \$500 in OpenAI API credits.
The ChatGPT interface was also equipped with several interactive features designed to collect user feedback, allowing users to give thumbs-up/thumbs-down (see Fig. 1) and open-text feedback about responses (see Fig. 2 and Fig. 3), as well as to regenerate a response and identify whether the second response was better, the same, or worse than the original (see Fig. 4). Additionally, checkboxes in the open-text feedback windows allowed users to structure their feedback against the prominent “helpful, honest, and harmless” framework for AI alignment (Askell et al., 2021). Other interfaces to LLMs, such as Anthropic’s Claude.ai and Google’s Bard/Gemini, incorporate feedback features similar to the ChatGPT interface to gather user input and preferences. The deployment of LLMs through these interfaces has made these systems accessible to a vast and diverse user base, presenting new challenges related to scale.
Figure 1. Thumbs-up or down feedback - ChatGPT (May 2023)
[…]
An image of the ChatGPT interface showing a prompt completion. To the right of ChatGPT’s response, there are two buttons side by side: a thumbs-up icon on the left and a thumbs-down icon on the right. This represents the simple binary feedback mechanism offered to users after each prompt completion.
Figure 2. Open-text window after thumbs-up - ChatGPT (May 2023)
[…]
An image of a pop-up window in the ChatGPT interface that appears after a user gives a thumbs-up rating. The window is titled “Provide additional feedback” and contains a large text box for users to input detailed feedback, in which the following text appears: “What do you like about the response?”. Below the text box to the right is a “Submit feedback” button.
Figure 3. Open-text window after thumbs-down - ChatGPT (May 2023)
[…]
An image of a pop-up window in the ChatGPT interface that appears after a user gives a thumbs-down rating. The window is titled “Provide additional feedback” and contains a large text box for users to input detailed feedback, in which the following text appears: “What was the issue with the response? How could it be improved?”. Below the text box are three checkboxes labelled “This is harmful/unsafe”, “This isn’t true”, and “This isn’t helpful”. At the bottom right of the window is a “Submit feedback” button.
Figure 4. Feedback after regenerating a response - ChatGPT (May 2023)
[…]
A screenshot of the ChatGPT interface showing a regenerated response. Between the responses is a set of three buttons arranged horizontally, allowing the user to rate whether the new response is “Better”, “Worse”, or the “Same” compared to the original response.
2.2.1. Scale and Affordances
The standardisation of feedback features across interfaces to different systems exemplifies the proceduralisation and replication that characterise other digital systems deployed at large scale (Gillespie, 2020). LLMs are based on deep neural networks that learn from massive datasets, and these systems have grown significantly in size in recent years—for example, from millions of parameters in GPT-1 to billions in GPT-2 (Solaiman et al., 2019) and GPT-3 (Brown et al., 2020), and (reportedly) trillions in GPT-4 (Patel and Wong, 2023). But more than simply expanding the size of a system, scaling digital systems across global contexts invariably requires imposing a standardised worldview and organising phenomena that can operate context-independently (Tsing, 2012; Hanna and Park, 2020). This sameness is achieved through massive infrastructural investments, such as data centres, standardised software and policies, and global distribution networks (Young et al., 2024). Feedback features in interfaces to LLMs (and the associated infrastructure for managing, storing and processing feedback data) represent one such infrastructural investment designed to elicit participation at scale. Notably, these feedback features are identical for all users of a given interface to an LLM (e.g., ChatGPT, Bard/Gemini, or Claude.ai) and similar across different interfaces to various LLMs, despite the diversity of global contexts in which users interact with these systems.
However, this sameness can obscure the power dynamics inherent in the feedback process. Achieving scale in digital contexts is an aggregation of power—engineering the suppression of context to render complex social realities coherent through a computationally legible frame (Tsing, 2012; Gillespie, 2020). As Eubanks (Eubanks, 2017) demonstrates, the design of scaled technical systems can reinforce existing inequalities and marginalise certain voices. For example, people less likely to engage with these interfaces, such as those with limited digital literacy or access, may be underrepresented in the feedback data used to refine the family of models underlying the interfaces. The standardisation of feedback features across different systems might also lead to the homogenisation of user input, risking the amplification of certain biases or the suppression of diverse perspectives. It is thus imperative to critically examine what kinds of perspectives and contexts these feedback features make visible and valuable as forms of participation—both to those who participate and to those who integrate feedback from participants into ML workflows. The concept of affordances provides one useful lens for conducting such an analysis.
2.2.2. Mechanisms and Conditions Framework of Affordances
Affordance theory states that an animal perceives its environment not only in terms of the shapes of objects or the space between objects but also in the possibilities of objects for action (Gibson, 1977). Animals, to which Gibson (Gibson, 1977) refers, share the same environment but perceive it differently according to their affordances—their capacity to benefit or experience harm from a particular feature of the environment. Norman (Norman, 2008) later adapted affordance theory for technology design, arguing that technological objects afford different possibilities for humans and other agents and that those possibilities cannot be defined except by reference to the object’s adoption and use (Hildebrandt, 2015). The affordances of a technological object are thus the link between the object’s features and the outcomes of the technological object’s adoption and use by subjects (Davis, 2020). Davis and Chouinard (Davis and Chouinard, 2016) describe these links as “relational processes among users, designers, environments, and things”.
Davis (Davis, 2020) operationalises the concept of affordances through a framework of mechanisms and conditions. Mechanisms describe the link between features of technological objects and the outcomes of their use by subjects, while conditions describe how affordances vary across different subjects and contexts. The framework proposes six mechanisms that contribute to the operationalisation of affordances: request, demand, encourage, discourage, refuse, and allow. These mechanisms are intended to serve as a tool for critical analysis rather than as a taxonomy of the features of technological objects, as each mechanism may reveal different aspects of a technical feature. Technological objects request and demand how user-subjects interact with them. Technological objects also encourage and discourage, refuse and allow actions of user-subjects. These mechanisms operate through three interrelated conditions: perception (how a subject perceives the artifact based on their skill and technical literacy), dexterity (what a subject can do with the artifact), and cultural and institutional legitimacy (the social and structural position of the user within institutions of power). The framework addresses binarism (assuming an artifact affords some function or not) and universalism (assuming an artifact operates identically for all) in affordance theory (Davis and Chouinard, 2016). Using this framework, we can explore “how, for whom, and under what circumstances” artefacts afford action (Davis, 2020).
If we recognise ML-enabled technologies as material artefacts that reflect and affect users through design decisions, and acknowledge their potential to exacerbate inequality in social settings and workplaces (Davis, 2023), we must scrutinise the affordances of such technologies and applications. The mechanisms and conditions framework provides a structured approach to analyse how the design of feedback features in interfaces to LLMs affords certain actions while constraining others, thereby shaping the nature and scope of user input and participation in RLHF workflows. Affordance theory is useful for analysing how users generate evaluative feedback through LLM interfaces and how this feedback is subsequently shaped and constrained to function as cybernetic input in RLHF workflows. The framework’s emphasis on the interdependent and power-laden relationship between humans and technologies supports our critical examination of the sameness of scaled systems. Furthermore, the framework offers a vocabulary for the (re)design of feedback features to address issues we may identify for user input and public participation through socio-technical analysis (Davis, 2020).
3. Method
We conducted an online survey to gather insights from ChatGPT users who had submitted feedback using the features in the browser-based interface. Through the survey, we aimed to explore feedback practices and perspectives of ChatGPT users to understand which users submit feedback, how they submit feedback, and why they submit feedback. We then drew on the insights from the survey to analyse the feedback features through the mechanisms and conditions framework of affordances. Below, we detail the process we used to conduct the survey and the analysis, and we then describe the limitations of our approach. In the following section, we present the insights from our analysis.
3.1. Data collection
3.1.1. Pilot Survey
We conducted two pilot surveys in March 2023 with 32 respondents from our university to refine our survey instrument’s clarity, flow, and length. In late March 2023, the Human Ethics Review Committee at our university approved our recruitment plan and final survey instrument.
3.1.2. Survey Instrument
The survey instrument consisted of 45 questions—approximately 21 were consistent across respondents, and the survey logic determined the remainder. The survey asked respondents about their frequency and types of ChatGPT usage, experiences with submitting feedback through the ChatGPT interface and external channels, reasons for submitting or not submitting feedback, perceptions of the feedback process, the likelihood of providing future feedback through the interface and external channels, and demographic characteristics (including the AI, ML, and data science background of respondents). Most questions were multiple-choice, with a few open-ended questions to gather more detailed insights and reflections from respondents.
3.1.3. Recruitment and Sampling
We administered the public survey on Qualtrics in May 2023 and used a multi-pronged approach to recruit participants. Initially, the authors posted the survey on their social media accounts (Twitter, Mastodon, and LinkedIn), forums frequented by active ChatGPT users (GitHub forums, Hugging Face forums, the OpenAI community forum, and Kaggle forums), and on the r/ChatGPT Discord server.
We also recruited participants through Prolific to address non-response bias and expand the respondent pool beyond those engaged with AI, ML, or data science in research or development capacities. We applied screening strategies on Prolific to invite relevant respondents to the final survey. Firstly, we used Prolific filters to identify respondents from any country, reporting fluency in English, an approval rate greater than or equal to 95%, and 10 or more previous submissions. Secondly, we invited filtered respondents to a custom screening survey with two questions asking participants about chatbot usage and feedback submission. [^1] ChatGPT usage was not an option for filters on Prolific in May 2023. Finally, we invited respondents approved after the screening survey to the final survey at three separate intervals to cover working hours in regions with different time zones. Respondents to the final survey on Prolific received the equivalent of £9/hr for their time.
3.1.4. Respondents
The survey received responses from 526 individuals—after removing incomplete surveys, failed attention checks [^2] Only Prolific surveys., and screened surveys [^3] Valid IP Address, “Yes” to “Read and agree to Participant Information Sheet”, “Yes” to “Has used ChatGPT”.—with a median completion time of five minutes and 11 seconds. The respondents were predominantly young, with over three quarters aged 18–34 (42.4% aged 18–24, 35.1% aged 25–34, n=524). The gender distribution was skewed towards men (63.1%), with 34.9% identifying as women, 1.2% as non-binary or third gender, and 0.8% preferring not to say or preferring to self-describe (n=521). Geographically, the respondents were primarily from Europe (55.4%), followed by Africa (15.6%), the United Kingdom (10.3%), and Oceania (6.5%) (n=507). The most common languages spoken by respondents were English (UK, US, or AU) (61.1%), Portuguese (13.4%), and Polish (12.2%) (n=509). [^4] Respondents could select up to three languages they speak daily.
Respondents reported varying levels of engagement with AI-related topics (e.g., through news articles, research papers, podcasts, etc.), with 63.1% at least weekly, 18.9% monthly, and only 18.0% rarely or never (n=523). A substantial proportion (41.0%) had used AI, ML, or data science tools beyond basic interaction or exploration before using ChatGPT (n=522). Furthermore, 26.5% had completed formal education or training in these areas (n=520), and 21.3% were working or had previously worked in a related field (n=521).
3.2. Data Analysis
Initially, we performed frequency analysis for most questions (see Appendix A). For certain variables measured on a 5-point Likert scale, we transformed the responses into three categories (e.g., Agree, Neutral, Disagree; or Likely, Neutral, Unlikely) to simplify the analysis and interpretation. We then performed chi-square tests of independence ($\chi^{2}$) to analyse relationships between responses, using Cramér’s V to determine the strength of the associations (Cohen, 2013). Before these analyses, we removed rows with missing values for demographics and experience factors and grouped categories with low frequencies for relevant questions (e.g., consolidating age groups and geographic regions). To address the multiple comparisons problem, we applied the Benjamini-Hochberg procedure to control the false discovery rate across all tests simultaneously (Benjamini and Hochberg, 1995). We report statistically significant results based on the adjusted $p$-values (adj. $p$-value). All statistical tests were conducted at a 95% confidence level (df=degrees of freedom; see Appendix B for full results).
For open-ended questions in the survey, we conducted a thematic analysis of responses (Braun and Clarke, 2006), with the first author initially developing codes based on the semantic content of the first 50 responses using tags in the coding software Atlas.ti. Both authors then met to discuss and consolidate the codes, after which the first author coded the remaining responses, followed by a final joint discussion to identify and agree on key themes reflecting the data. Finally, we drew on insights from the survey and our interface analysis to inform the affordances analysis.
3.3. Limitations
Our methodology has several important limitations. Firstly, we recognise that our survey dataset does not statistically represent any group—neither ChatGPT users in general nor the subset of ChatGPT users who have submitted feedback through the interface (Tufekci, 2014). We sampled a subset of ChatGPT users in May 2023, though the networks of the authors and the population of users on Prolific influence our sample. The number and characteristics of ChatGPT users in general are not freely available, nor are the number or characteristics of ChatGPT users who have submitted feedback. This survey instrument was conducted to sample that population within a constrained budget. As such, our survey insights are exploratory and not generalisable to the population of ChatGPT users. Nevertheless, these insights are a useful starting point for understanding user practices and perspectives related to feedback features in interfaces to LLMs.
Secondly, our survey relies on self-reported data from respondents. We did not create a primary dataset of feedback instances, as we could not access a dataset of feedback instances submitted through the ChatGPT interface. Instead, we collected a secondary dataset of the type of feedback (e.g., thumbs-up/down, open-text feedback, etc.) and the broad categories of feedback (e.g., correcting technical errors/bugs, social and ethical information, etc.) provided by respondents. There may be discrepancies between what users reported and their actual feedback behaviour. We note, however, that the survey was conducted primarily to understand the practices and perspectives of ChatGPT users who have submitted feedback rather than to analyse the content of the feedback itself.
We also engaged with early adopters of ChatGPT, as our survey was conducted in the first six months after deployment. In April and May 2023, surveys of U.S. adults found that only 16% (Vogels, 2023) and 14% (Berndt et al., 2023) had used ChatGPT, respectively. As a result, the respondents’ demographics and responses to questions about their experience with AI, ML, and data science may not represent the broader population of ChatGPT users as adoption changes over time. Early adopters tend to have higher education levels, higher socioeconomic status, and be more engaged with developments in emerging technologies than later adopters (Rogers, 2003), which may skew the survey results. If ChatGPT usage becomes more mainstream, the characteristics and perspectives of users may change, potentially affecting the types of feedback provided and the motivations for engaging with the feedback features. However, as of March 2024, only 23% of U.S. adults reported using ChatGPT (McClain, 2024), suggesting that the early adopter effect may still be relevant to our findings.
Finally, while survey findings provide useful information about user practices and perspectives concerning feedback, there are limitations to using this data for our affordances analysis. The survey provides a snapshot of user perspectives and experiences at a point in time. The ChatGPT interface and feedback features have and will continue to change, so the affordances of those feedback features and user behaviours may change accordingly. However, as of May 2024, the basic functionality has not changed significantly, though users are prompted with feedback categories where they were previously provided with an open-text box (see Fig. 5), and a user may be requested to submit pairwise feedback (see, e.g., Fig. 6). We also note that user interviews or observations could provide more in-depth information about how users perceive and interact with the feedback features, though we argue that survey data is sufficient for our high-level analysis through the mechanisms and conditions framework of affordances.
Figure 5. Additional feedback categories after selecting thumbs-down - ChatGPT (May 2024)
[…]
An image of the ChatGPT interface showing a feedback window that appears after a user selects the thumbs-down option. The window is titled “Provide additional feedback” and presents multiple categorised feedback options for users to specify why they disliked the response. Categories include “Don’t like the style”, “Not factually correct”, “Didn’t follow instructions”, “Refused when it shouldn’t have”, “Being lazy”, “Unsafe or problematic”, and “Other”. At the bottom of the window is a single-line text box labelled “(Optional) Feel free to add specific details” for users to provide more specific feedback.
Figure 6. Pairwise feedback options - ChatGPT (May 2024, content of responses hidden)
[…]
An image of the ChatGPT interface displaying a pairwise comparison feature. At the top of the window, the text reads “Which response do you prefer? Your answer will help make ChatGPT better”. Below, the image shows two text areas side by side containing different generated responses, though the content of the responses is hidden for privacy. This interface design facilitates direct comparison between two alternative responses.
4. Insights and Analysis
This section reports insights from the survey regarding ChatGPT usage, feedback practices, and respondents’ motivations for submitting feedback. Throughout the section, we draw on the insights to inform our analysis of the affordances of feedback features in the ChatGPT interface.
Our analysis suggests how the design of feedback features imposes constraints on end-user participation. In summary, the interface requests feedback formatted in the ‘helpful, honest, and harmless’ framework, demands interaction before allowing feedback submission, and encourages simple, frequent, individualised, and unidirectional feedback, particularly thumbs-up/down ratings and feedback after regenerating a response. However, the interface discourages collective feedback and refuses feedback modification over time. Additionally, our survey indicates that the feedback features may appeal more to users with frequent engagement, prior experience, and knowledge of AI, ML or data science.
4.1. Mechanisms of Affordance
Our survey reveals how users interact with ChatGPT and its feedback features, which influence the type and quality of data fed back into RLHF workflows. Respondents used ChatGPT frequently and in diverse ways. Over 71.6% (n=525) used ChatGPT at least weekly, and the most common use cases were answering specific questions (e.g., factual, technical, or theoretical questions), generating text (e.g., drafting emails, writing essays or reports), and academic research assistance (e.g., generating citations, summarising articles).
This usage pattern is significant because a person cannot directly influence ML workflows outside that system. A user must prompt the dialogue agent before they can submit feedback. This design feature effectively demands interaction with an LLM before allowing participation through feedback submission, establishing the first key mechanism of affordance we observed.
4.1.1. Submitting Feedback
The feedback features in the ChatGPT interface are designed to elicit individual responses, which effectively discourages collective feedback or discussions among users. This individualisation extends to the permanence of feedback: users cannot modify their input once submitted, even if they gain new information or their perspective changes. While users could theoretically input the same prompt a second time and submit feedback, the stochastic nature of LLMs means each response is likely to be unique, making direct comparisons or revisions challenging. This design choice reinforces the individual nature of the feedback process and limits opportunities for evolving, collective input.
Our survey findings indicate that the ChatGPT interface encourages thumbs-up/down feedback and feedback after regenerating a response. Most respondents (84.2%, n=526) indicated they had submitted feedback through one of the three methods in the ChatGPT interface. Thumbs-up/down feedback was the most common (76.0%, n=526), followed by feedback after regenerating responses (62.9%, n=526), while open-text feedback was less common (29.0%, n=526).
Respondents submitting thumbs-up/down feedback also reported doing so most frequently (53.3% every, most, or about half the times they had used ChatGPT, n=400), followed by feedback after regenerating a response (47.4% every, most, or about half the times they had used ChatGPT, n=331). Respondents submitting open-text feedback reported doing so least frequently (26.4% every, most, or about half the times they had used ChatGPT, n=152).
Respondents who submitted open-text feedback (n=152) were primarily motivated by improving the performance and capabilities of the system. We asked this group of respondents to reflect on the most recent time they submitted open-text feedback and report their motivations for doing so. Respondents could select multiple motivations, but those who selected multiple motivations were forced to choose their primary motivation in the following question. Taking the weighted frequencies, the primary motivations were (in order) to: “contribute to AI development”, “comment on use-case specific feedback”, “report a technical issue or bug”, and “suggest new features or improvements” (see Table 12 in Appendix A). Expressing social or ethical concerns about system outputs was the least common motivation (excluding the range of individual responses in the “other” option).
4.1.2. Reflecting on Feedback
We also asked respondents who submitted open-text feedback what feedback they thought was most useful for improving ChatGPT’s performance or impact. The responses revealed a hierarchy of feedback themes. The most prominent themes were technical issues (bugs, glitches, incorrect information) and system performance (efficiency, speed, and quality of responses). Less common but still significant were concerns about accuracy (‘hallucinations’ or false information) and the reliability and truthfulness of responses. Though less frequent again, some users highlighted social and ethical concerns (including harmful, biased, or misleading responses). Finally, some feedback focused on suggesting new features or improvements.
Most respondents were ambivalent about whether their reasons for submitting feedback had been addressed—47.4% neither agreed nor disagreed, 23.7% somewhat agreed, and 16.4% somewhat disagreed (n=152). Few of those who submitted open-text feedback indicated they had received follow-up communication or updates from OpenAI (11.2%, n=152), though the vast majority of those who had indicated they were satisfied with the response or resolution.
The feedback features in the ChatGPT interface allow for open-text feedback, including commentary on social and ethical issues. However, our survey results reveal that users predominantly focus on performance-related concerns when submitting feedback. This focus appears to be shaped by the affordances of the feedback system itself rather than by explicitly discouraging social or ethical feedback. Despite OpenAI’s initial blog post requesting feedback on risks and harms, users who responded to our survey seem to have interpreted the purpose of their participation primarily through the lens of improving ChatGPT’s performance. This pattern highlights how the design of feedback features can effectively constrain user perspectives and shape the nature of user input.
4.2. Conditions
4.2.1. Perception
Respondents generally perceived the feedback features in the interface as accessible and user-friendly. Over three-quarters reported confidence in their ability to evaluate and submit feedback about the performance and/or impact of generative AI systems such as ChatGPT (“Strongly Agree” (21.8%) or “Somewhat Agree” (56.5%), n=524). Of those who submitted at least one of the three feedback modes, almost all found doing so through the features extremely or somewhat easy (95.1% n=442).
However, some users were more confident than others in submitting feedback. Younger users (18-34) and those who frequently engaged with AI-related topics expressed more confidence than would be expected if there was no relationship between these factors and confidence. Conversely, users with less frequent engagement were less confident than would be expected. We found weak associations between confidence and age ($\chi^{2}$ = 17.068, adj. $p$-value = 0.045, df = 6, Cramér’s V = 0.131) and frequent engagement with AI-related topics ($\chi^{2}$2 = 20.726, adj. $p$-value = 0.045, df = 8, Cramér’s V = 0.144).
4.2.2. Dexterity
Few respondents reported barriers to accessing ChatGPT or submitting feedback, suggesting that users with varying levels of dexterity can provide input. When asked about barriers to accessing ChatGPT or submitting feedback, most respondents (77.3%, n=519) reported experiencing no barriers. Among those who encountered barriers, the most common were language and internet connectivity issues. We did not find significant associations between most demographic factors or AI experience and knowledge factors and those who experienced barriers. While those working in AI were more likely to report experiencing barriers than would be expected if there was no relationship between the variables, the association was weak ($\chi^{2}$ = 9.821, adj. $p$-value = 0.023, df = 1, Cramér’s V = 0.143).
However, the requirement to prompt the dialogue agent before submitting feedback means only users can submit feedback through the interface’s features. Our sample, therefore, does not include non-user perspectives, and given that most respondents used ChatGPT at least weekly, it also includes limited input from infrequent users. Notably, of those who completed the survey but did not indicate they had submitted feedback through one of the three options, the most common reason was that they were unaware of the feedback option (34.6%, n=78).
While the ChatGPT interface may constrain direct feedback, respondents found alternative ways to express their opinions and participate in discussions about the system. Most respondents (80.4%, n=526) did not indicate they shared feedback about ChatGPT through other channels or platforms. However, those who did primarily used X/Twitter (42%), followed by Facebook (18%) and Reddit (17%). Several respondents engaged with dedicated communities, including subreddits like r/ChatGPT and online forums like the OpenAI Developer Forum. Most respondents who did indicate they had shared feedback through other channels or platforms (71.9%, n=103) reported they were likely to continue to share feedback through those channels.
4.2.3. Cultural and Institutional Legitimacy
Interactive feedback features may appeal more to users with certain cultural and institutional backgrounds, which can influence who submits feedback and the types of feedback they provide. Indeed, most of our survey respondents were young and identified as men, which reflects ChatGPT’s reported user base (Vogels, 2023) and other studies of contributors to human feedback datasets (Kirk et al., 2024). Generally, feedback types were proportional across demographic groups, though we found that respondents’ region had a moderate association with providing open-text feedback ($\chi^{2}$ = 13.990, adj. $p$-value = 0.023, df = 3, Cramér’s V = 0.192).
Instead, our survey results suggest that feedback features appeal more to users with frequent engagement or familiarity with AI, ML, and data science. Frequent ChatGPT users, particularly daily users, were more likely to provide thumbs-up/down feedback ($\chi^{2}$ = 13.689, adj. $p$-value = 0.045, df = 4, Cramér’s V = 0.166) than would be expected if there was no relationship between the variables, and those who previously used AI tools ($\chi^{2}$ = 8.926, adj. $p$-value = 0.023 df = 1, Cramér’s V = 0.154) were more likely to provide open-text feedback than would be expected. We also observed moderate associations suggesting more frequent ChatGPT users and those often engaging with AI-related topics were more likely to submit open-text feedback than would be expected if there was no relationship between these variables. However, these associations did not reach statistical significance after adjusting $p$-values. Additionally, we noted a borderline significant association between education or training in AI and open-text feedback submission (adj. $p$-value = 0.059).
4.3. Summary of Affordances
Table 1 outlines the affordances of the feedback features in the ChatGPT interface based on the analysis above, organised by the six mechanisms proposed by the framework. The affordances are linked to the conditions (perception, dexterity, and cultural and institutional legitimacy) that shape how users interact with and perceive the feedback features.
Table 1. Summary of Affordances
Mechanism Affordance Condition
Request Feedback formatted in the ‘helpful, honest, and harmless’ framework Feedback from users to developers (Dexterity)
Demand Interaction with ChatGPT before feedback submission Status as a user (Dexterity)
Encourage Thumbs-up/down feedback; Feedback after regenerating a response; Performance-related open-text feedback Ease of use and accessibility of interface (Perception); Frequency of use and familiarity with the system (Perception, Dexterity); AI/ML experience and knowledge (Cultural and institutional legitimacy)
Discourage Feedback from collectives; Discussions among users Individualised feedback design (Cultural and institutional legitimacy)
Refuse Modifying feedback over time Stochastic nature of the system (Dexterity)
Allow Alternative channels for feedback and discussion (e.g., social media, online forums) Availability of alternative channels (Dexterity); Accessibility and appeal of feedback features to certain user groups (Cultural and institutional legitimacy)
5. Discussion
Participation is never fully open or anarchic; it is always formatted with “specific rules and scripts” (Kelty, 2020). Feedback features in interfaces to LLMs format user input into RLHF workflows. Therefore, the design of these features shapes the legitimacy of user participation in the iteration of LLMs. Our analysis of feedback features in ChatGPT suggests that the feedback format imposed by the interface is simple, individualised, unidirectional, and focused on performance-related concerns. This format significantly constrains user participation and risks becoming self-perpetuating—users may internalise limited roles in the iteration of LLMs, discouraging them from seeking broader engagement or viewing themselves as co-creators in LLM development. This phenomenon is not unique to LLMs; similar patterns have been observed in other technological domains, such as social media platforms (Gillespie, 2020).
Affordance theory as a design tool requires an orientation that “technologies, and the worlds they build, need not be as they are and can, instead, be otherwise” (Davis, 2023). As LLMs increasingly become infrastructure, it is important to consider participation beyond mere preference elicitation. In this section, we discuss two directions for the (re)design of feedback processes. Initially, we review recent developments in deliberative approaches to feedback, such as recent democratic experiments of OpenAI and Anthropic. However, we argue that users and stakeholders can never get to the heart of closed systems by commenting on model outputs. Instead, we encourage a shift in approach from alignment of model outputs to specific user preferences to dialogue with the public about the purpose and applications of LLMs, and attention to the ongoing infrastructuring work required to support and sustain diverse ‘publics’ over time.
5.1. Deliberative Processes
As previously noted, the existing feedback features request feedback from users to developers and discourage both feedback from groups of users and discussion amongst users. In other words, feedback through existing features is atomised and unidirectional (Wang et al., 2023). Additionally, as the number of ChatGPT users has grown and the system has stabilised over time, user input has been constrained from a wide to a narrow aperture—from free-text feedback forms to open-text feedback windows to singular buttons focused on crowdsourcing inputs to RLHF. These latest feedback formats are inadequate for resolving social and ethical considerations, requiring us to consider other design approaches.
If algorithmic fairness is primarily a political task, as Wong (Wong, 2020) argues, rather than solely a technical task, we must consider how to resolve issues of bias and fairness politically rather than technically. Alongside aggregating citizen preferences through voting, deliberation among citizens is a key element of democratic legitimacy (Fishkin, 1991; Knight and Johnson, 1994; Curato et al., 2017). Deliberative democracy involves an association of members who govern their affairs by deliberation among the members (Elster and Przeworski, 1998)—making collective decisions determined by those subject to the decisions: not only their preferences but also their reasoning (Cohen, 2005). Such associations tend to be between broad groups of people—in this context, users, developers, and potentially groups of both—and are directed towards participants considering others’ perspectives and potentially changing their own over time.
Interestingly, survey respondents identified several other platforms on which they discuss their experiences and preferences regarding ChatGPT, engaging in the kind of collective deliberation that the existing feedback features discourage. For example, OpenAI discussion forums allow communication between users and developers and subreddits such as r/ChatGPT allow discussion amongst users (without monitoring or oversight by OpenAI). However, it is unclear whether or how these discussions are input into development processes; while they function as evaluative feedback, there is no clear cybernetic feedback loop.
Some technology companies have been experimenting with deliberative modes for feedback about AI systems. Recent efforts by OpenAI to fund research into technical solutions for “democratic inputs into AI” (Eloundou and Lee, 2024) and experiments by Anthropic to draft a constitution for Claude through deliberative processes (Huang et al., 2024) have attracted significant attention. These efforts go beyond eliciting and aggregating user preferences towards encouraging deliberation amongst people with diverse perspectives (including non-users) and demonstrating how the preferences and reasoning of multiple participants are input (as cybernetic feedback) into development workflows. In addition, Meta has experimented with Community Forums to deliberate on the rulesets for Generative AI (Broxmeyer, 2024), and Google has developed the STELA process for deliberative discussions with historically underrepresented groups in the U.S.A. to understand their priorities and concerns with respect to model outputs (Bergman et al., 2024).
While these processes address some of the limitations of existing feedback features, they remain largely experimental and peripheral to the dominant method of feedback submission through browser-based interfaces. Most of these approaches (with the Google example as an exception) take as a given that consensus on the output of the process is desirable and achievable. However, refusal, agonism and dissensus are key elements of democratic processes (DiSalvo, 2015, 2010; Zong and Matias, 2024). When consulting multiple and diverse communities, values and preferences for AI design may be in conflict (Sambasivan et al., 2021; Qadri et al., 2023) and difficult to reconcile (Santurkar et al., 2023; Durmus et al., 2023; Miller et al., 2007). This is particularly the case where the processes are managed by those with commercial interests in consensus outcomes, reflecting long-standing concerns in participatory design (Sanders and Stappers, 2014).
Moreover, such democratic inputs to the alignment problem, much like the existing feedback processes they are intended to augment or replace, narrowly scope participation around the output of specific families of models rather than engaging participants in earlier stages of the development process. Although deliberative processes may address some limitations of existing mechanisms and change how companies like OpenAI and Anthropic make decisions about model behaviour, they do not significantly alter the power dynamics between users, the public, and platforms. Users and other participants remain as consumers rather than co-creators, limited in their influence over core aspects of earlier stages of development.
5.2. Infrastructuring Participation
Our study reveals a tension between technology companies’ stated intentions and their implementation of feedback features in LLMs. ChatGPT exemplifies this disparity: what OpenAI initially sought (diverse, substantive feedback to uncover potential risks and harms) gave way to what they could effectively use within their existing workflows (easily processed, formatted data). This change not only constrains user input but also, as our survey indicates, favours frequent ChatGPT users and those with AI experience, narrowing the group of users influencing LLM iteration.
While eliciting user preferences as input into RLHF workflows may be well-intentioned and useful for technology companies to improve model performance, it is inadequate for the larger task facing us all. We all face the challenge of publicly, deliberatively, and inclusively addressing the ongoing tectonic shifts in information and knowledge generation across economic, social, cultural, and political domains before LLMs become unremarkable and unquestionable infrastructure.
If we are to shift our focus from aligning model outputs to specific user preferences towards dialogue with the public about the purpose and applications of LLMs, then we must think critically about who will form this ‘public’, how it will form, and what channels facilitate its participation. Currently, technology companies configure participation from atomised users of dialogue agents like ChatGPT, tightly defining processes for feedback on model outputs. In ChatGPT’s case, the ‘public’ is constructed primarily to address a matter of concern to OpenAI rather than those engaged as participants. Moreover, groups or institutions affected by the development and deployment of LLMs lack direct mechanisms to influence these feedback processes (Wang et al., 2023).
Technology companies might argue for increasing the number and diversity of participants providing feedback through existing channels. However, this approach misses the core issue. Drawing parallels with Arnstein’s (Arnstein, 1969) critique of participation in the context of urban planning, we argue that identifying problems in the proceduralisation of participation through feedback features in LLMs is not a call for more involvement but for a fundamental rethinking of the process. We echo Wang et al. (Wang et al., 2023) in urging companies to focus on the participation of groups or institutions representing populations likely to be significantly impacted by a given LLM. Infrastructuring such participation requires deep engagement with the methodological tools of Participatory Design (PD), which has a long-standing practice of critical thinking about how to form and sustain ‘publics’ to address participants’ concerns over time (DiSalvo, 2009; Le Dantec and DiSalvo, 2013). When implemented effectively, PD prioritises future issues concerning participants rather than solely focusing on the immediate concerns of designers or researchers.
Recent research offers promising directions for rethinking processes and infrastructure for participation in LLM development. The STELA process (Bergman et al., 2024) demonstrates one method for conducting community-centred deliberative discussions about LLMs, albeit focused on aligning models with the preferences of specific groups of people. Suresh et al. (Suresh et al., 2024) expand on this concept by proposing a comprehensive framework for domain-specific community infrastructure, including shared technical resources, norms, and governance structures. Their approach aims to enable more meaningful participation by groups and institutions in shaping the development and application of foundation models in specific domains. Building on these ideas, we encourage participatory workshops with stakeholders in fields like education, healthcare, and journalism as a practical mechanism for realising and refining such domain-specific infrastructure. However, implementing these participatory approaches faces significant challenges: maintaining local relevance while extending across domains and regions, translating insights into actionable development inputs, and securing sustainable funding for ongoing infrastructure in the rapidly changing landscape of AI development. To advance meaningful public participation in the iteration of LLMs, the focus of research and effort must be on addressing these challenges rather than defaulting to procedural and perfunctory engagement.
6. Conclusion
Our analysis reveals that existing feedback features in the ChatGPT interface constrain participation around developers’ instrumental needs. This approach to user input and participation views engagement as a means to an end rather than an opportunity for dialogue. While deliberative approaches may address some limitations of existing mechanisms, the scope of user and non-user input remains constrained. To move beyond perfunctory engagement, we call for a shift in focus from user preference elicitation to building robust infrastructures for sustained, meaningful dialogue about the purpose and application of LLMs. As these technologies become increasingly embedded in information, workplace, and entertainment services, the urgency of this task cannot be overstated.
Appendix A Survey Results
Appendix A compiles the survey results from our study. It contains a series of tables with detailed breakdowns of respondent usage patterns, feedback behaviours, demographics, and perceptions of submitting feedback using the interactive features in the ChatGPT interface.
Table 2. Frequency of ChatGPT Usage
Frequency Count Percentage
A few times a week 216 41.1%
Once a week 96 18.3%
Monthly 84 16.0%
Rarely 65 12.4%
Daily 64 12.2%
Total 525 100.0%
(Response rate: 525/526)
Table 3. Ways of Using ChatGPT
Usage Count Percentage
Answering specific questions 375 71.3%
Generating text 332 62.9%
Academic research assistance 268 51.0%
Entertainment or casual conversation 226 43.0%
Professional or technical problem-solving 222 42.2%
Creative writing 214 40.7%
Language learning or translation 122 23.3%
Personal advice or emotional support 109 20.7%
Other 22 4.2%
(Response rate: 526/526, Multiple selections allowed)
Table 4. Submitted Thumbs-Up/Down Feedback
Submitted feedback Count Percentage
Yes 400 76.0%
No 126 24.0%
Total 526 100.0%
(Response rate: 526/526)
Table 5. Frequency of Thumbs-Up/Down Feedback
Frequency Count Percentage
Sometimes I have used ChatGPT 139 34.8%
About half the time I have used ChatGPT 98 24.5%
Most times I have used ChatGPT 92 23.0%
Rarely 48 12.0%
Every time I have used ChatGPT 23 5.8%
Total 400 100.0%
(Response rate: 400/400)
Table 6. Submitted Open-Text Feedback
Submitted feedback Count Percentage
No 248 47.1%
Yes 152 29.0%
N/A 126 23.9%
Total 526 100.0%
(Response rate: 526/526, N/A = “No” to submitting thumbs-up/down feedback.)
Table 7. Submitted Open-Text Feedback (“Yes” to thumbs-up/down feedback)
Submitted feedback Count Percentage
No 248 62.0%
Yes 152 38.0%
Total 400 100.0%
(Response rate: 400/400)
Table 8. Frequency of Open-Text Feedback
Frequency Count Percentage
Sometimes I have used ChatGPT 77 50.7%
Rarely 35 23.0%
Most times I have used ChatGPT 20 13.2%
About half the time I have used ChatGPT 17 11.2%
Every time I have used ChatGPT 3 2.0%
Total 152 100.0%
(Response rate: 152/152)
Table 9. Submitted Feedback After Regenerating a Response
Submitted feedback Count Percentage
Yes 331 62.9%
No 195 37.1%
Total 526 100.0%
(Response rate: 526/526)
Table 10. Frequency of Feedback After Regenerating a Response
Frequency Count Percentage
Sometimes I have regenerated a response 120 36.3%
Most times I have regenerated a response 67 20.2%
About half the time I have regenerated a response 65 19.6%
Rarely 54 16.3%
Every time I have regenerated a response 25 7.6%
Total 331 100.0%
(Response rate: 331/331)
Table 11. Reasons for Not Submitting Feedback
Reasons Count Percentage
Unaware of feedback option 27 34.6%
Not enough time or interest in submitting feedback 21 26.9%
Satisfied with the performance and have not felt the need to submit feedback 17 21.8%
Unsure how to provide feedback 7 9.0%
Other 5 6.4%
Unsatisfied with the performance, but do not believe feedback will lead to improvements 1 1.3%
Total 78 100.0%
(Response rate: 78/78)
Table 12. Motivations for Providing Feedback
Motivation Weighted frequency
To contribute to AI development 159
To comment on use-case specific feedback 94
To report a technical issue or bug 74
To suggest new features or improvements 73
To express social or ethical concerns about ChatGPT 32
Other 15
(Response rate: 152/152)
Note: Weighted frequencies based on a two-part question. Part A (multiple selection) responses weighted as 1; Part B (main motivation) responses weighted as 2.
Table 13. Received Follow-up Regarding Feedback
Response Count Percentage
No 135 88.8%
Yes 17 11.2%
Total 152 100.0%
(Response rate: 152/152)
Table 14. Satisfaction with Follow-up
Satisfaction level Count Percentage
Somewhat satisfied 10 58.8%
Extremely satisfied 5 29.4%
Neither satisfied nor dissatisfied 2 11.8%
Total 17 100.0%
(Response rate: 17/17)
Table 15. Reasons for Submitting Feedback Addressed
Agreement level Count Percentage
Neither agree nor disagree 72 47.4%
Somewhat agree 36 23.7%
Somewhat disagree 25 16.4%
Strongly disagree 11 7.2%
Strongly agree 8 5.3%
Total 152 100.0%
(Response rate: 152/152)
Table 16. Ease of Submitting Feedback
Level of ease Count Percentage
Extremely easy 292 66.1%
Somewhat easy 128 29.0%
Neither easy nor difficult 18 4.1%
Somewhat difficult 4 0.9%
Total 442 100.0%
(Response rate: 442/443)
Table 17. Likelihood of Submitting Feedback in Future
Likelihood Count Percentage
Somewhat likely 255 48.7%
Extremely likely 161 30.7%
Neither likely nor unlikely 67 12.8%
Somewhat unlikely 34 6.5%
Extremely unlikely 7 1.3%
Total 524 100.0%
(Response rate: 524/526)
Table 18. Provided Feedback via Other Platforms or Channels
Provided feedback Count Percentage
No 423 80.4%
Yes 103 19.6%
Total 526 100.0%
(Response rate: 526/526)
Table 19. Other Channels or Platforms
Channels Count Percentage
X/Twitter 42 40.8%
Facebook 18 17.5%
Other 18 17.5%
Reddit forums 17 16.5%
Discord servers 17 16.5%
LinkedIn 10 9.7%
Contacting OpenAI (e.g., via email) 10 9.7%
Other online discussion forums 9 8.7%
Mastodon 3 2.9%
(Response rate 103/103, Multiple selections allowed)
Table 20. Likelihood of Submitting Feedback via Other Channels or Platforms in Future
Likelihood Count Percentage
Somewhat likely 46 44.7%
Extremely likely 28 27.2%
Neither likely nor unlikely 15 14.6%
Somewhat unlikely 10 9.7%
Extremely unlikely 4 3.9%
Total 103 100.0%
(Response rate: 103/103)
Table 21. Age
Age Count Percentage
18-24 222 42.4%
25-34 184 35.1%
35-44 72 13.7%
45-54 31 5.9%
55-64 8 1.5%
65+ 4 0.8%
Prefer not to say 3 0.6%
Total 524 100.0%
(Response rate: 524/526)
Table 22. Gender
Gender Count Percentage
Man 329 63.1%
Woman 182 34.9%
Non-binary / third gender 6 1.2%
Prefer not to say 3 0.6%
Prefer to self-describe 1 0.2%
Total 521 100.0%
(Response rate: 521/526)
Table 23. Region
Region Count Percentage
Europe 281 55.4%
Africa 79 15.6%
United Kingdom 52 10.3%
Oceania 33 6.5%
North America 23 4.5%
Central America 21 4.1%
South America 9 1.8%
Asia 9 1.8%
Total 507 100.0%
(Response rate: 507/526)
Table 24. Languages Spoken Daily (Grouped)
Languages Count Percentage
English (UK) 163 32.0%
English (US) 116 22.8%
Portuguese 68 13.4%
Polish 62 12.2%
Italian 33 6.5%
English (AU) 32 6.3%
Zulu 29 5.7%
Spanish (including Latin American) 49 9.6%
Greek 15 2.9%
German (Standard and Swiss) 14 2.8%
Sotho 13 2.6%
French (including Canadian) 9 1.8%
Other African languages 8 1.6%
Dutch 6 1.2%
Czech 5 1.0%
Other European languages 14 2.8%
Other Asian languages 11 2.2%
Other languages 7 1.4%
Total 509 100.0%
(Response rate: 509/526, up to 3 selections allowed)
Notes: ‘Other African languages’ includes Setswana, Xhosa, Yoruba, Swahili, and Hausa; ’Other European languages’ includes Finnish, Hungarian, Latvian, Russian, Norwegian, Danish, Swedish, Catalan, Ukrainian, and Lithuanian; ’Other Asian languages’ includes Hindi, Hebrew, Turkish, Arabic, Mongolian, Korean, Japanese, Persian, Bengali, Tagalog, Thai, and Sinhalese; ’Other languages’ includes Portuguese (Brazilian), Slovenian, Croatian, and unspecified others.
Table 25. Barriers Accessing ChatGPT or Submitting Feedback
Barriers Count Percentage
No barriers 401 77.3%
Barriers 118 22.7%
Total 519 100.0%
(Response rate: 519/526)
Table 26. Frequency of Engaging with AI-related Topics
Frequency Count Percentage
Weekly 243 46.5%
Monthly 99 18.9%
Daily 87 16.6%
Rarely 85 16.3%
Never 9 1.7%
Total 523 100.0%
(Response rate: 523/526)
Table 27. Prior Use of AI, ML, or Data Science Tools
Response Count Percentage
No 308 59.0%
Yes 214 41.0%
Total 522 100.0%
(Response rate: 522/526)
Table 28. Completed Education or Training in AI, ML, or Data Science
Response Count Percentage
No 382 73.5%
Yes 138 26.5%
Total 520 100.0%
(Response rate: 520/526)
Table 29. Currently or Previously Working in AI, ML, or Data Science
Response Count Percentage
No 410 78.7%
Yes 111 21.3%
Total 521 100.0%
(Response rate: 521/526)
Table 30. Confidence Submitting Feedback
Confidence level Count Percentage
Somewhat agree 296 56.5%
Strongly agree 114 21.8%
Neither agree nor disagree 79 15.1%
Somewhat disagree 30 5.7%
Strongly disagree 5 1.0%
Total 524 100.0%
(Response rate: 524/526)
Appendix B Statistical Relationships and Test Results
Appendix B presents our statistical analysis of relationships between key variables in our study. Table 31 presents the results of chi-square tests of associations between various factors such as user characteristics, feedback behaviours, and perceptions of submitting feedback through the ChatGPT interface.
Table 31. Chi-square Test Results
Relationship $\chi^{2}$ $p$ Adj. $p$ df CV Interp. $n$
Thumbs-Up/down & Regenerating feedback 50.419 ¡.001 ¡.001* 1 .319 Mod. 497
Education/Training in AI & Confidence 12.773 .002 .023* 2 .160 Weak 497
Working in AI & Barriers 9.821 .002 .023* 1 .143 Weak 477
Region & Open-text feedback 13.990 .003 .023* 3 .192 Mod. 378
Prior AI use & Open-text feedback 8.926 .003 .023* 1 .154 Weak 378
Thumbs-Up/down & ChatGPT frequency 13.689 .008 .045* 4 .166 Weak 497
AI topics frequency & Confidence 20.726 .008 .045* 8 .144 Weak 497
Age & Confidence 17.068 .009 .045* 6 .131 Weak 497
Education/Training in AI & Open-text 6.132 .013 .059 1 .127 Weak 378
Gender & Regenerating feedback 10.427 .015 .061 3 .145 Weak 497
Age & Barriers 9.372 .025 .078 3 .140 Weak 477
Open-text & ChatGPT frequency 10.956 .027 .078 4 .170 Mod. 378
AI topics frequency & Open-text 10.937 .027 .078 4 .170 Mod. 378
Region & Regenerating feedback 9.151 .027 .078 3 .136 Weak 497
Working in AI & Open-text feedback 4.660 .031 .080 1 .111 Weak 378
Age & Regenerating feedback 8.678 .034 .080 3 .132 Weak 497
Prior AI use & Barriers 4.512 .034 .080 1 .097 Negl. 477
Education/Training in AI & Barriers 4.010 .045 .101 1 .092 Negl. 477
Gender & Confidence 12.513 .051 .103 6 .112 Weak 497
Prior AI use & Confidence 5.953 .051 .103 2 .109 Weak 497
Working in AI & Confidence 5.632 .060 .114 2 .106 Weak 497
Open-text & Regenerating feedback 3.424 .064 .117 1 .095 Negl. 378
AI topics frequency & Thumbs-Up/down 7.615 .107 .186 4 .124 Weak 497
Working in AI & Regenerating feedback 2.375 .123 .205 1 .069 Negl. 497
Regenerating & ChatGPT frequency 5.004 .287 .459 4 .100 Weak 497
Age & Open-text feedback 3.342 .342 .526 3 .094 Weak 378
Gender & Thumbs-Up/down 3.160 .368 .545 3 .080 Weak 497
AI topics frequency & Barriers 3.962 .411 .554 4 .091 Weak 477
AI topics frequency & Regenerating 3.932 .415 .554 4 .089 Weak 497
Prior AI use & Thumbs-Up/down 0.666 .415 .554 1 .037 Negl. 497
Gender & Barriers 2.482 .478 .607 3 .072 Weak 477
Education/Training in AI & Thumbs 0.468 .494 .607 1 .031 Negl. 497
Working in AI & Thumbs-Up/down 0.453 .501 .607 1 .030 Negl. 497
Gender & Open-text feedback 1.801 .615 .723 3 .069 Weak 378
Age & Thumbs-Up/down 1.713 .634 .725 3 .059 Negl. 497
Prior AI use & Regenerating feedback 0.140 .708 .773 1 .017 Negl. 497
Region & Barriers 1.361 .715 .773 3 .053 Negl. 477
Education/Training in AI & Regenerating 0.065 .799 .841 1 .011 Negl. 497
Region & Thumbs-Up/down 0.708 .871 .894 3 .038 Negl. 497
Region & Confidence 1.548 .956 .956 6 .039 Negl. 497
Notes:
  • Adj. $p$ = $p$-value after Benjamini-Hochberg procedure; df = Degrees of Freedom; CV = Cramér’s V; Interp. = Interpretation; Mod. = Moderate; Negl. = Negligible, $n$ = Sample size after removing missing values.
  • Table organised in order of adj. $p$-value.
  • * Indicates statistically significant $p$-value after adjustment.
  • $\chi^{2}$ test was not performed between Thumbs-Up/down & Open-text feedback. Open-text feedback was only available to users who provided Thumbs-Up/down feedback, making these variables dependent.
  • Table 32 provides a guide for interpreting Cramér’s V values, in accordance with (Cohen, 2013).
Table 32. Interpretation of Cramér’s V
Association df = 1 df = 2 df $\geq$ 3
Weak 0.10 - ¡0.30 0.07 - ¡0.21 0.06 - ¡0.17
Moderate 0.30 - ¡0.50 0.21 - ¡0.35 0.17 - ¡0.29
Strong $\geq$0.50 $\geq$0.35 $\geq$0.30

### v2

@@
 Large language models (LLMs) are now accessible to anyone with a computer, a web browser, and an internet connection via browser-based interfaces, shifting the dynamics of participation in AI development.
-This paper examines the affordances of interactive feedback features in ChatGPT’s interface, analysing how they shape user input and participation in LLM iteration.
-Drawing on a survey of ChatGPT users and applying the mechanisms and conditions framework of affordances, we demonstrate that these features encourage simple, frequent, and performance-focused feedback while discouraging collective input and discussions among users.
-We argue that this feedback format significantly constrains user participation, reinforcing power imbalances between users, the public, and companies developing LLMs.
-Our analysis contributes to the growing body of literature on participatory AI by critically examining the limitations of existing feedback processes and proposing directions for their redesign.
-To enable more meaningful public participation in AI development, we advocate for a shift away from processes focused on aligning model outputs with specific user preferences.
-Instead, we emphasise the need for processes that facilitate dialogue between companies and diverse ‘publics’ about the purpose and applications of LLMs.
-This approach requires attention to the ongoing work of infrastructuring—creating and sustaining the social, technical, and institutional structures necessary to address matters of concern to groups impacted by AI development and deployment.
+This article examines how interactive feedback features in ChatGPT’s interface afford user participation in LLM iteration.
+Drawing on a survey of early ChatGPT users and applying the mechanisms and conditions framework of affordances, we analyse how these features shape user input.
+Our analysis indicates that these features encourage simple, frequent, and performance-focused feedback while discouraging collective input and discussions among users.
+Drawing on participatory design literature, we argue such constraints, if replicated across broader user bases, risk reinforcing power imbalances between users, the public, and companies developing LLMs.
+Our analysis contributes to the growing literature on participatory AI by critically examining the limitations of existing feedback processes and proposing directions for redesign.
+Rather than focusing solely on aligning model outputs with specific user preferences, we advocate for creating infrastructure that supports sustained dialogue about the purpose and applications of LLMs.
+This approach requires attention to the ongoing work of “infrastructuring”—creating and sustaining the social, technical, and institutional structures necessary to address matters of concern to stakeholders impacted by LLM development and deployment.
 Large language models, human-AI interaction, user feedback, participatory AI
-copyright: none
+journal: JRCjournalvolume: 2journalnumber: 4article: 16publicationmonth: 10journalyear: 2025doi: 10.1145/3748516ccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Artificial intelligenceccs: Social and professional topics User characteristicsccs: Applied computing Sociology;
 1. Introduction
 The deployment of ChatGPT by OpenAI in 2022 marked a significant shift in how large language models (LLMs) are made available to the public.
-ChatGPT’s browser-based interface allows broad access to interactions with LLMs through text prompts, expanding user participation in the iterative development of LLMs.
+ChatGPT’s browser-based interface facilitates broad access to interactions with LLMs through text prompts, expanding user participation in the iterative development of LLMs.
 Previously, companies such as OpenAI and Google took a more cautious approach.
@@
   1. (1)
-    What are the affordances of interactive feedback features in the ChatGPT interface?
+    What feedback features are present in the ChatGPT interface, and how are they used?
   2. (2)
-    How do those affordances shape the nature and scope of user input and participation in the iteration of LLMs?
+    How do these feedback features, through specific mechanisms and conditions, afford particular types of user participation in the iteration of LLMs?
 To address these questions, we first surveyed ChatGPT users who have submitted feedback.
-We then drew on the survey insights to analyse interactive features in the ChatGPT interface that allow users to submit feedback, using Davis’s (Davis, 2020) mechanisms and conditions framework of affordances.
-Our analysis indicates that the ChatGPT interface encourages simple, frequent, individualised, and unidirectional feedback on performance-related concerns while discouraging feedback from collectives or discussions among users.
-We discuss implications for (re)designing feedback processes for LLMs.
-First, we explore deliberative approaches that might enable bidirectional communication, changes in perspectives over time, and discussion about social or ethical issues among groups of users.
-However, we recognise that deliberation on rulesets for model outputs will not fundamentally alter the power dynamics between users, the public, and platforms.
-Second, we consider the “infrastructuring” (Le Dantec and DiSalvo, 2013) required to support and sustain more expansive feedback channels between the public and technology companies.
-We conclude by urging companies deploying LLMs to shift their focus from aligning model outputs with user preferences to fostering sustained dialogue with diverse ‘publics’ about the purposes and applications of these technologies.
-This approach requires careful attention to the ongoing work of infrastructuring participation—creating and maintaining the social, technical, and institutional structures necessary for genuine public engagement in AI development.
-We must move beyond preference elicitation before LLMs become ubiquitous infrastructure—systems that are embedded in other structures, transparent in use, and only visible upon breakdown (Star and Ruhleder, 1996).
+We then drew on the survey insights to analyse feedback features in the ChatGPT interface that allow users to submit feedback, using Davis’s (Davis, 2020) mechanisms and conditions framework of affordances.
+Understanding early adopter use of these feedback features offers insights into their initial affordances and potential limitations when scaling to broader populations; these early patterns are likely to shape the norms and standards for participation in LLM iteration.
+Our analysis indicates that the ChatGPT interface encourages simple, frequent, individualised, and unidirectional feedback on performance-related concerns while discouraging feedback from collectives or discussions among users, and refuses feedback modification over time.
+Further, technical background and engagement frequency significantly influence participation patterns.
+We discuss implications arising from the constraints on user participation observed in our study for (re)designing feedback for LLMs.
+First, we explore deliberative approaches that might enable bidirectional communication between users and developers, changes in perspectives over time, and discussion about social or ethical issues among groups of users.
+Second, we consider the “infrastructuring” (Le Dantec and DiSalvo, 2013) that could support more expansive feedback channels, to enable sustained dialogue with stakeholders affected by LLM deployments about the purposes and applications of these technologies.
+Moving beyond simple preference elicitation toward more expansive feedback channels appears increasingly important as LLMs become ubiquitous and unquestionable infrastructure—systems embedded in other structures, transparent in use, and only visible upon breakdown (Star and Ruhleder, 1996).
 2. Background
@@
 Interest in participatory approaches for aligning LLMs is part of a broader participatory turn in AI research, design, and development (Delgado et al., 2023).
-Participatory Al projects involve end users and those impacted by AI systems in the design and development of those systems (Cooper and Zafiroglu, 2024), drawing on long-standing traditions in other fields, such as participatory design (Muller and Kuhn, 1993; Sundblad, 2011) and participatory planning (Arnstein, 1969; Kelty, 2020).
-The turn to participation in AI research, design and development has included papers, panels and workshops on participatory machine learning (ML) at prominent computer science conferences (Kulynych et al., 2020; Zytko et al., 2022) and calls for public participation in AI in recent policy frameworks (Farthing et al., 2021; U.S.
+Participatory AI projects involve end users and those impacted by AI systems in the design and development of those systems (Cooper and Zafiroglu, 2024), drawing on long-standing traditions in other fields, such as participatory design (Muller and Kuhn, 1993; Sundblad, 2011) and participatory planning (Arnstein, 1969; Kelty, 2020).
+Participatory AI projects now feature in articles, panels and workshops at prominent computer science conferences (Kulynych et al., 2020; Zytko et al., 2022) and in recent policy frameworks (Farthing et al., 2021; U.S.
 Office of Science and Technology Policy, 2022; Tabassi, 2023).
+The participatory turn is one response to calls for moves away from formal models of fairness (Green, 2022) toward situated collaborations with communities affected by AI systems (Sloane, 2022).
+These calls reflect a growing recognition that algorithmic fairness is not merely a technical problem amenable to formal solutions, but a socio-political challenge requiring attention to structural inequities (Hoffmann, 2019; Wong, 2020).
+Indeed, technical approaches alone may obscure or reinforce structural inequities, leading some to advocate for alternatives to algorithmic fairness such as “algorithmic reparation” (Davis et al., 2021).
 Participatory approaches promise more “democratic” AI systems (Kulynych et al., 2020) and more “pluralistic” alignment (Kirk et al., 2024).
 However, there is a broad spectrum of participatory practices in AI design, from informing or consulting participants to equitable partnerships with participants (Birhane et al., 2022; Corbett et al., 2023; Delgado et al., 2023).
-This spectrum of participatory practices also existed in participatory design, participatory planning, and other antecedent fields.
 For the development of large-scale AI systems, participation has often taken passive and involuntary forms through the representation of groups of individuals in datasets used to train models (Denton et al., 2020; Sloane et al., 2022).
-While feedback features in interfaces to LLMs provide a more active and voluntary form of user participation, their emergence as the primary mode for the public to influence development processes has raised questions about the conceptual leap from structured preference elicitation to genuine participatory engagement (Feffer et al., 2023b).
-Additionally, there is a significant gap between the scale of participatory methods and the training and deployment of recent AI systems (Diaz and Madaio, 2023).
+While feedback features in interfaces to LLMs provide a more active and voluntary form of user participation, their emergence as the primary mode for publics to influence development processes has raised a question about the conceptual leap from structured preference elicitation to genuine participatory engagement (Feffer et al., 2023b).
+Additionally, there is a significant gap between the scale of participatory methods in computing research and the scale at which recent AI systems are trained and deployed (Diaz and Madaio, 2025).
 Surveys of participatory approaches in computing research (Cooper et al., 2022) and participatory AI projects (Delgado et al., 2023) indicate that engagements have primarily been localised, focusing on one or a few groups or communities with shared characteristics.
@@
 The concept of feedback loops influenced subsequent AI research directions (Minsky, 1961) and remains fundamental to AI systems that learn over time.
-Concurrently, ‘feedback’ has also become a general term referring to information provided by an observer (e.g., users or other stakeholders) assessing a system’s performance.
-Reinforcement Learning from Human Feedback (RLHF) demonstrates how the two types of feedback are combined in modern AI development.
-RLHF begins with evaluative feedback, when a model generates two responses to a given prompt and human raters determine which response is better than the other according to criteria.
+Concurrently, “feedback” has also become a general term referring to information provided by an observer (e.g., users or other stakeholders) assessing a system’s performance.
+Reinforcement Learning from Human Feedback (RLHF) combines these feedback types in modern AI development.
+RLHF begins with evaluative feedback, when a model generates two responses to a given prompt and human raters determine which response is better than the other according to some criteria.
 These human assessments are then used to create a cybernetic feedback loop: an RL algorithm uses the assessments to predict how a human would rate other responses, and an LLM is fine-tuned on the resulting reward model (Ziegler et al., 2019; Ouyang et al., 2022; Bai et al., 2022).
-This approach has become central to aligning pre-trained LLMs with specific preferences (Lambert et al., 2023).
-However, integrating human feedback into the learning process increases the time to train a model and the labour costs involved (Maslej et al., 2024).
-RLHF requires large, diverse groups of human raters to assess prompt completions before fine-tuning the model in the final stage (Prabhakaran et al., 2021).
-The importance of RLHF for aligning LLMs has increased the demand for human feedback data (Lambert et al., 2023) and research attention to practices for creating and curating diverse, representative, and quality human feedback data (Feffer et al., 2023a; Kirk et al., 2024; Prabhakaran et al., 2021).
-However, less attention has been paid to human feedback collected through interactive features in interfaces to LLMs, in which users evaluate prompt completions of deployed systems.
-This form of feedback is captured after a given LLM has been trained and deployed, and such feedback may be used to improve future model iterations.
+While central to aligning pre-trained LLMs with specific preferences (Lambert et al., 2023), integrating human feedback into the learning process increases training time and labour costs (Maslej et al., 2024).
+RLHF requires large, diverse groups of human raters to assess prompt completions before fine-tuning the model (Prabhakaran et al., 2021).
+The growing importance of RLHF has increased demand for human feedback data (Lambert et al., 2023) and research attention on practices for creating and curating diverse, representative, and quality human feedback data (Feffer et al., 2023a; Kirk et al., 2024; Prabhakaran et al., 2021).
+However, less attention has been paid to human feedback collected through interfaces to LLMs, in which users evaluate prompt completions of deployed systems.
+This post-deployment feedback may be used to improve future model iterations.
 For example, OpenAI used feedback on non-factual responses obtained through the feedback features in ChatGPT’s interface, with additional labelled comparison data, to train reward models that reduced GPT-4’s tendency to “hallucinate” compared to GPT-3.5 (OpenAI, 2023).
@@
 Users were encouraged to submit free-form feedback via a Google form that could help OpenAI identify potential real-world risks and harms, with winners receiving \$500 in OpenAI API credits.
-The ChatGPT interface was also equipped with several interactive features designed to collect user feedback, allowing users to give thumbs-up/thumbs-down (see Fig.
-1) and open-text feedback about responses (see Fig.
-2 and Fig.
-3), as well as to regenerate a response and identify whether the second response was better, the same, or worse than the original (see Fig.
-4).
-Additionally, checkboxes in the open-text feedback windows allowed users to structure their feedback against the prominent “helpful, honest, and harmless” framework for AI alignment (Askell et al., 2021).
-Other interfaces to LLMs, such as Anthropic’s Claude.ai and Google’s Bard/Gemini, incorporate feedback features similar to the ChatGPT interface to gather user input and preferences.
+Figure 1. Thumbs-up or down feedback - ChatGPT (May 2023).:
+Figure 2. Open-text window after thumbs-up - ChatGPT (May 2023).:
+The ChatGPT interface was also equipped with several interactive features, including: thumbs-up/thumbs-down ratings (see Figure 1) open-text feedback (see Figure 2 and Figure 3), and comparative feedback after a user regenerates a response (see Figure 4).
+Checkboxes in the open-text feedback windows allow users to structure their feedback against the prominent “helpful, honest, and harmless” framework for AI alignment (Askell et al., 2021).
+Similar features appear in other interfaces to LLMs, such as Google’s Bard/Gemini and Anthropic’s Claude.ai.
 The deployment of LLMs through these interfaces has made these systems accessible to a vast and diverse user base, presenting new challenges related to scale.
-Figure 1. Thumbs-up or down feedback - ChatGPT (May 2023)
-[…]
-An image of the ChatGPT interface showing a prompt completion.
-To the right of ChatGPT’s response, there are two buttons side by side: a thumbs-up icon on the left and a thumbs-down icon on the right.
-This represents the simple binary feedback mechanism offered to users after each prompt completion.
-Figure 2. Open-text window after thumbs-up - ChatGPT (May 2023)
-[…]
-An image of a pop-up window in the ChatGPT interface that appears after a user gives a thumbs-up rating.
-The window is titled “Provide additional feedback” and contains a large text box for users to input detailed feedback, in which the following text appears: “What do you like about the response?”.
-Below the text box to the right is a “Submit feedback” button.
-Figure 3. Open-text window after thumbs-down - ChatGPT (May 2023)
-[…]
-An image of a pop-up window in the ChatGPT interface that appears after a user gives a thumbs-down rating.
-The window is titled “Provide additional feedback” and contains a large text box for users to input detailed feedback, in which the following text appears: “What was the issue with the response?
-How could it be improved?”.
-Below the text box are three checkboxes labelled “This is harmful/unsafe”, “This isn’t true”, and “This isn’t helpful”.
-At the bottom right of the window is a “Submit feedback” button.
-Figure 4. Feedback after regenerating a response - ChatGPT (May 2023)
-[…]
-A screenshot of the ChatGPT interface showing a regenerated response.
-Between the responses is a set of three buttons arranged horizontally, allowing the user to rate whether the new response is “Better”, “Worse”, or the “Same” compared to the original response.
+Figure 3. Open-text window after thumbs-down - ChatGPT (May 2023).:
+Figure 4. Feedback after regenerating a response - ChatGPT (May 2023).:
 2.2.1. Scale and Affordances
 The standardisation of feedback features across interfaces to different systems exemplifies the proceduralisation and replication that characterise other digital systems deployed at large scale (Gillespie, 2020).
-LLMs are based on deep neural networks that learn from massive datasets, and these systems have grown significantly in size in recent years—for example, from millions of parameters in GPT-1 to billions in GPT-2 (Solaiman et al., 2019) and GPT-3 (Brown et al., 2020), and (reportedly) trillions in GPT-4 (Patel and Wong, 2023).
+LLMs have grown exponentially in size in recent years—from billions of parameters in GPT-2 (Solaiman et al., 2019) to (reportedly) trillions in GPT-4 (Patel and Wong, 2023).
 But more than simply expanding the size of a system, scaling digital systems across global contexts invariably requires imposing a standardised worldview and organising phenomena that can operate context-independently (Tsing, 2012; Hanna and Park, 2020).
 This sameness is achieved through massive infrastructural investments, such as data centres, standardised software and policies, and global distribution networks (Young et al., 2024).
-Feedback features in interfaces to LLMs (and the associated infrastructure for managing, storing and processing feedback data) represent one such infrastructural investment designed to elicit participation at scale.
+Feedback features (and the associated infrastructure for managing, storing, and processing feedback data) represent one such infrastructural investment designed to elicit participation at scale.
 Notably, these feedback features are identical for all users of a given interface to an LLM (e.g., ChatGPT, Bard/Gemini, or Claude.ai) and similar across different interfaces to various LLMs, despite the diversity of global contexts in which users interact with these systems.
@@
 The standardisation of feedback features across different systems might also lead to the homogenisation of user input, risking the amplification of certain biases or the suppression of diverse perspectives.
+The “helpful, honest, and harmless” framework (Askell et al., 2021), for example, is commonly used to structure feedback for RLHF workflows.
+However, applying such standardised frameworks risks imposing particular values across diverse contexts (Gabriel and Keeling, 2025).
+Definitions of “helpfulness” can vary culturally (Sambasivan et al., 2021), and “honesty” privileges certain epistemic standards inherent in a model’s training data (Denton et al., 2020), while determining “harmlessness” involves complex and contested judgments about potentialharms (Weidinger et al., 2022; Hoffmann, 2019; Gabriel and Keeling, 2025).
 It is thus imperative to critically examine what kinds of perspectives and contexts these feedback features make visible and valuable as forms of participation—both to those who participate and to those who integrate feedback from participants into ML workflows.
@@
 Affordance theory states that an animal perceives its environment not only in terms of the shapes of objects or the space between objects but also in the possibilities of objects for action (Gibson, 1977).
-Animals, to which Gibson (Gibson, 1977) refers, share the same environment but perceive it differently according to their affordances—their capacity to benefit or experience harm from a particular feature of the environment.
+Animals share the same environment but perceive it differently according to their affordances—their capacity to benefit or experience harm from a particular feature of the environment.
 Norman (Norman, 2008) later adapted affordance theory for technology design, arguing that technological objects afford different possibilities for humans and other agents and that those possibilities cannot be defined except by reference to the object’s adoption and use (Hildebrandt, 2015).
 The affordances of a technological object are thus the link between the object’s features and the outcomes of the technological object’s adoption and use by subjects (Davis, 2020).
-Davis and Chouinard (Davis and Chouinard, 2016) describe these links as “relational processes among users, designers, environments, and things”.
-Davis (Davis, 2020) operationalises the concept of affordances through a framework of mechanisms and conditions.
+Davis and Chouinard (2016) describe these links as “relational processes among users, designers, environments, and things”.
+Davis (2020) operationalises the concept of affordances through a framework of mechanisms and conditions.
+The framework extends affordance theory by reframing the fundamental question away from what technologies afford to “how technologies afford, for whom, and under what circumstances(Davis, 2020).
 Mechanisms describe the link between features of technological objects and the outcomes of their use by subjects, while conditions describe how affordances vary across different subjects and contexts.
-The framework proposes six mechanisms that contribute to the operationalisation of affordances: request, demand, encourage, discourage, refuse, and allow.
+The framework proposes six mechanisms that contribute to the operationalisation of affordances: request, demand, encourage, discourage, refuse, and allow (Davis, 2020).
 These mechanisms are intended to serve as a tool for critical analysis rather than as a taxonomy of the features of technological objects, as each mechanism may reveal different aspects of a technical feature.
 Technological objects request and demand how user-subjects interact with them.
+Requests refer to bids placed by the artifact initiating some action or set of actions, while demands represent more forceful bids that leave users with little room for non-compliance.
 Technological objects also encourage and discourage, refuse and allow actions of user-subjects.
-These mechanisms operate through three interrelated conditions: perception (how a subject perceives the artifact based on their skill and technical literacy), dexterity (what a subject can do with the artifact), and cultural and institutional legitimacy (the social and structural position of the user within institutions of power).
-The framework addresses binarism (assuming an artifact affords some function or not) and universalism (assuming an artifact operates identically for all) in affordance theory (Davis and Chouinard, 2016).
-Using this framework, we can explore “how, for whom, and under what circumstances” artefacts afford action (Davis, 2020).
+These latter mechanisms represent responses from the artifact to activities that users seek to enact.
+Allow functions neutrally, applying to bids placed by both the artifact and user-subjects engaging with it.
+Importantly, these mechanisms are not mutually exclusive but operate simultaneously—features that request certain feedback formats may encourage some user behaviour at the same time as discouraging others.
+The mechanisms operate through three interrelated conditions: perception (how a subject perceives the artifact based on their skill and technical literacy), dexterity (what a subject can do with the artifact), and cultural and institutional legitimacy (the social and structural position of the user within institutions of power) (Davis, 2020).
+The interaction between mechanisms and conditions creates variations in how technologies afford actions for different users.
+For example, without sufficient dexterity or the requisite cultural and institutional legitimacy, what may be allowed becomes discouraged or refused (Davis, 2023).
+The framework thus addresses binarism (assuming an artifact affords some function or not) and universalism (assuming an artifact operates identically for all) in affordance theory (Davis and Chouinard, 2016).
 If we recognise ML-enabled technologies as material artefacts that reflect and affect users through design decisions, and acknowledge their potential to exacerbate inequality in social settings and workplaces (Davis, 2023), we must scrutinise the affordances of such technologies and applications.
 The mechanisms and conditions framework provides a structured approach to analyse how the design of feedback features in interfaces to LLMs affords certain actions while constraining others, thereby shaping the nature and scope of user input and participation in RLHF workflows.
-Affordance theory is useful for analysing how users generate evaluative feedback through LLM interfaces and how this feedback is subsequently shaped and constrained to function as cybernetic input in RLHF workflows.
-The framework’s emphasis on the interdependent and power-laden relationship between humans and technologies supports our critical examination of the sameness of scaled systems.
 Furthermore, the framework offers a vocabulary for the (re)design of feedback features to address issues we may identify for user input and public participation through socio-technical analysis (Davis, 2020).
@@
 We conducted an online survey to gather insights from ChatGPT users who had submitted feedback using the features in the browser-based interface.
-Through the survey, we aimed to explore feedback practices and perspectives of ChatGPT users to understand which users submit feedback, how they submit feedback, and why they submit feedback.
+Through the survey, we aimed at exploring feedback practices and perspectives of ChatGPT users to understand which users submit feedback, how they submit feedback, and why they submit feedback.
 We then drew on the insights from the survey to analyse the feedback features through the mechanisms and conditions framework of affordances.
@@
 3.1.1. Pilot Survey
-We conducted two pilot surveys in March 2023 with 32 respondents from our university to refine our survey instrument’s clarity, flow, and length.
-In late March 2023, the Human Ethics Review Committee at our university approved our recruitment plan and final survey instrument.
+We conducted two pilot surveys in March 2023 with 32 respondents from the Australian National University (ANU) to refine our survey instrument’s clarity, flow, and length.
+In late March 2023, the ANU approved our recruitment plan and final survey instrument.
 3.1.2. Survey Instrument
-The survey instrument consisted of 45 questions—approximately 21 were consistent across respondents, and the survey logic determined the remainder.
-The survey asked respondents about their frequency and types of ChatGPT usage, experiences with submitting feedback through the ChatGPT interface and external channels, reasons for submitting or not submitting feedback, perceptions of the feedback process, the likelihood of providing future feedback through the interface and external channels, and demographic characteristics (including the AI, ML, and data science background of respondents).
-Most questions were multiple-choice, with a few open-ended questions to gather more detailed insights and reflections from respondents.
+The survey instrument consisted of 36 substantive questions exploring respondents’ frequency and types of ChatGPT usage, experiences submitting feedback through the ChatGPT interface and external channels, reasons for submitting or not submitting feedback, perceptions of the feedback process, the likelihood of providing future feedback, and demographic characteristics (see Appendix A for survey questions).
+Most questions were multiple-choice, with several open-ended questions included to gather more detailed insights and reflections.
 3.1.3. Recruitment and Sampling
-We administered the public survey on Qualtrics in May 2023 and used a multi-pronged approach to recruit participants.
-Initially, the authors posted the survey on their social media accounts (Twitter, Mastodon, and LinkedIn), forums frequented by active ChatGPT users (GitHub forums, Hugging Face forums, the OpenAI community forum, and Kaggle forums), and on the r/ChatGPT Discord server.
-We also recruited participants through Prolific to address non-response bias and expand the respondent pool beyond those engaged with AI, ML, or data science in research or development capacities.
-We applied screening strategies on Prolific to invite relevant respondents to the final survey.
+We administered the public survey on Qualtrics in May 2023 and used a multi-pronged approach to recruit a diverse range of participants.
+This approach was necessary as the ChatGPT user population was not well-defined at the time of our study.
+Initially, we posted the survey on our social media accounts (X (formerly known as Twitter), Mastodon, and LinkedIn), forums frequented by active ChatGPT users (GitHub forums, Hugging Face forums, the OpenAI community forum, and Kaggle forums), and on the r/ChatGPT Discord server.
+We then expanded recruitment through Prolific to address potential non-response bias and broaden the respondent pool.
+We applied screening strategies on Prolific to invite suitable respondents to the final survey.
 Firstly, we used Prolific filters to identify respondents from any country, reporting fluency in English, an approval rate greater than or equal to 95%, and 10 or more previous submissions.
@@
 3.1.4. Respondents
-The survey received responses from 526 individuals—after removing incomplete surveys, failed attention checks [^2] Only Prolific surveys., and screened surveys [^3] Valid IP Address, “Yes” to “Read and agree to Participant Information Sheet”, “Yes” to “Has used ChatGPT”.—with a median completion time of five minutes and 11 seconds.
-The respondents were predominantly young, with over three quarters aged 18–34 (42.4% aged 18–24, 35.1% aged 25–34, n=524).
-The gender distribution was skewed towards men (63.1%), with 34.9% identifying as women, 1.2% as non-binary or third gender, and 0.8% preferring not to say or preferring to self-describe (n=521).
-Geographically, the respondents were primarily from Europe (55.4%), followed by Africa (15.6%), the United Kingdom (10.3%), and Oceania (6.5%) (n=507).
-The most common languages spoken by respondents were English (UK, US, or AU) (61.1%), Portuguese (13.4%), and Polish (12.2%) (n=509).
+The survey received responses from 526 individuals—after removing incomplete surveys, failed attention checks, [^2] Only Prolific surveys.
+and screened surveys [^3] Valid IP Address, “Yes” to “Read and agree to Participant Information Sheet”, “Yes” to “Has used ChatGPT”.—with a median completion time of five minutes and 11 seconds.
+We recruited 82 respondents through the first recruitment step and 445 through the second.
+As shown in Table 1, our sample includes perspectives from users with varying levels of technical experience and engagement with AI-related topics (e.g., through news articles, research papers, podcasts).
+Most respondents did not have formal AI training (73.5%) or relevant work experience (78.7%), suggesting our recruitment approach reached beyond technically experienced users.
+However, respondents were predominantly young, with over three quarters aged 18–34, and the gender distribution was skewed toward men.
+Geographically, the majority of respondents were from Europe, and the most common languages spoken by respondents were English (UK, US, or AU), Portuguese, and Polish.
 [^4] Respondents could select up to three languages they speak daily.
-Respondents reported varying levels of engagement with AI-related topics (e.g., through news articles, research papers, podcasts, etc.), with 63.1% at least weekly, 18.9% monthly, and only 18.0% rarely or never (n=523).
-A substantial proportion (41.0%) had used AI, ML, or data science tools beyond basic interaction or exploration before using ChatGPT (n=522).
-Furthermore, 26.5% had completed formal education or training in these areas (n=520), and 21.3% were working or had previously worked in a related field (n=521).
+Table 1. Respondent Demographics
+Characteristic Distribution
+Age ($n$ = 524) 18-34 (77.5%), 35-54 (19.6%), 55+ (2.3%)
+Gender ($n$ = 521) Man (63.1%), Woman (34.9%), Non-binary/Third gender (1.2%), Prefer not to say/Prefer to self-describe (0.8%)
+Region ($n$ = 507) Europe (55.4%), Africa (15.6%), UK (10.3%), Oceania (6.5%), Other (12.2%)
+AI Experience Prior use of AI/ML tools (41.0%, $n$ = 522), Formal education in AI/ML (26.5%, $n$ = 520), Working in AI/ML field (21.3%, $n$ = 521)
+AI Engagement ($n$ = 523) Daily/Weekly (63.1%), Monthly (18.9%), Rarely/Never (18.0%)
 3.2. Data Analysis
-Initially, we performed frequency analysis for most questions (see Appendix A).
-For certain variables measured on a 5-point Likert scale, we transformed the responses into three categories (e.g., Agree, Neutral, Disagree; or Likely, Neutral, Unlikely) to simplify the analysis and interpretation.
-We then performed chi-square tests of independence ($\chi^{2}$) to analyse relationships between responses, using Cramér’s V to determine the strength of the associations (Cohen, 2013).
+Initially, we performed frequency analysis for most questions (see Tables 937).
+For certain variables measured on a five-point Likert scale, we transformed the responses into three categories (e.g., Agree, Neutral, Disagree; or Likely, Neutral, Unlikely) to enhance statistical validity for chi-squared tests (by increasing cell counts and reducing the risk of minimum frequency violations) and improve interpretability.
+To assess the potential impact of this categorisation on our ordinal Likert scale data we performed sensitivity analyses on key relationships using the original five-point scales.
+These analyses showed consistent significance patterns in most cases, with similar effect sizes across both approaches.
+We then performed chi-squared tests of independence ($\chi^{2}$) to analyse relationships between responses, using Cramér’s V to determine the strength of the associations (Cohen, 2013).
 Before these analyses, we removed rows with missing values for demographics and experience factors and grouped categories with low frequencies for relevant questions (e.g., consolidating age groups and geographic regions).
@@
 $p$-value).
-All statistical tests were conducted at a 95% confidence level (df=degrees of freedom; see Appendix B for full results).
+All statistical tests were conducted at a 95% confidence level (df=degrees of freedom; see Appendix C for full results).
 For open-ended questions in the survey, we conducted a thematic analysis of responses (Braun and Clarke, 2006), with the first author initially developing codes based on the semantic content of the first 50 responses using tags in the coding software Atlas.ti.
@@
 Our methodology has several important limitations.
-Firstly, we recognise that our survey dataset does not statistically represent any group—neither ChatGPT users in general nor the subset of ChatGPT users who have submitted feedback through the interface (Tufekci, 2014).
-We sampled a subset of ChatGPT users in May 2023, though the networks of the authors and the population of users on Prolific influence our sample.
-The number and characteristics of ChatGPT users in general are not freely available, nor are the number or characteristics of ChatGPT users who have submitted feedback.
-This survey instrument was conducted to sample that population within a constrained budget.
-As such, our survey insights are exploratory and not generalisable to the population of ChatGPT users.
-Nevertheless, these insights are a useful starting point for understanding user practices and perspectives related to feedback features in interfaces to LLMs.
-Secondly, our survey relies on self-reported data from respondents.
+Firstly, the survey dataset does not statistically represent all ChatGPT users or the subset who submit feedback, limiting the generalisability of our findings (Tufekci, 2014).
+We sampled a subset of users in May 2023, influenced by the authors’ networks and the Prolific user population.
+Additionally, while necessary to engage with users during ChatGPT’s early release, sourcing some respondents from dedicated online forums might introduce a bias toward users inclined to discuss experiences publicly, potentially overlapping with those who share feedback through alternative channels.
+Our multi-pronged recruitment approach was pragmatic given that characteristics of the user base and feedback-submitting population were undefined at the time of study.
+Consequently, our survey insights are exploratory but useful as a starting point for understanding user practices and perspectives related to feedback features.
+Additionally, our survey relies on self-reported data from respondents.
 We did not create a primary dataset of feedback instances, as we could not access a dataset of feedback instances submitted through the ChatGPT interface.
-Instead, we collected a secondary dataset of the type of feedback (e.g., thumbs-up/down, open-text feedback, etc.) and the broad categories of feedback (e.g., correcting technical errors/bugs, social and ethical information, etc.) provided by respondents.
+Instead, we collected a secondary dataset of the type of feedback (e.g., thumbs-up/down, open-text feedback) and the broad categories of feedback (e.g., correcting technical errors/bugs, social and ethical information) provided by respondents.
 There may be discrepancies between what users reported and their actual feedback behaviour.
@@
 The ChatGPT interface and feedback features have and will continue to change, so the affordances of those feedback features and user behaviours may change accordingly.
-However, as of May 2024, the basic functionality has not changed significantly, though users are prompted with feedback categories where they were previously provided with an open-text box (see Fig.
-5), and a user may be requested to submit pairwise feedback (see, e.g., Fig.
-6).
+However, as of May 2024, the basic functionality has not changed significantly, though users are prompted with feedback categories where they were previously provided with an open-text box (see Figure 5), and a user may be requested to submit pairwise feedback (see, e.g., Figure 6).
 We also note that user interviews or observations could provide more in-depth information about how users perceive and interact with the feedback features, though we argue that survey data is sufficient for our high-level analysis through the mechanisms and conditions framework of affordances.
-Figure 5. Additional feedback categories after selecting thumbs-down - ChatGPT (May 2024)
-[…]
-An image of the ChatGPT interface showing a feedback window that appears after a user selects the thumbs-down option.
-The window is titled “Provide additional feedback” and presents multiple categorised feedback options for users to specify why they disliked the response.
-Categories include “Don’t like the style”, “Not factually correct”, “Didn’t follow instructions”, “Refused when it shouldn’t have”, “Being lazy”, “Unsafe or problematic”, and “Other”.
-At the bottom of the window is a single-line text box labelled “(Optional) Feel free to add specific details” for users to provide more specific feedback.
-Figure 6. Pairwise feedback options - ChatGPT (May 2024, content of responses hidden)
-[…]
-An image of the ChatGPT interface displaying a pairwise comparison feature.
-At the top of the window, the text reads “Which response do you prefer?
-Your answer will help make ChatGPT better”.
-Below, the image shows two text areas side by side containing different generated responses, though the content of the responses is hidden for privacy.
-This interface design facilitates direct comparison between two alternative responses.
+Figure 5. Additional feedback categories after selecting thumbs-down - ChatGPT (May 2024).:
+Figure 6. Pairwise feedback options - ChatGPT (May 2024, content of responses hidden).:
 4. Insights and Analysis
 This section reports insights from the survey regarding ChatGPT usage, feedback practices, and respondents’ motivations for submitting feedback.
-Throughout the section, we draw on the insights to inform our analysis of the affordances of feedback features in the ChatGPT interface.
-Our analysis suggests how the design of feedback features imposes constraints on end-user participation.
-In summary, the interface requests feedback formatted in the ‘helpful, honest, and harmless’ framework, demands interaction before allowing feedback submission, and encourages simple, frequent, individualised, and unidirectional feedback, particularly thumbs-up/down ratings and feedback after regenerating a response.
-However, the interface discourages collective feedback and refuses feedback modification over time.
-Additionally, our survey indicates that the feedback features may appeal more to users with frequent engagement, prior experience, and knowledge of AI, ML or data science.
+Our analysis indicates how the design of feedback features imposes constraints on end-user participation, operating through various mechanisms and conditions that we explore below.
+First, we present findings about user interactions with ChatGPT, then examine each affordance mechanism, and finally discuss how these mechanisms operate through different conditions.
+Respondents used ChatGPT frequently and in diverse ways.
+Over 71.6% (nn=525) used ChatGPT at least weekly, and the most common use cases were answering specific questions (e.g., factual, technical, or theoretical questions), generating text (e.g., drafting e-mails, writing essays or reports), and academic research assistance (e.g., generating citations, summarising articles).
 4.1. Mechanisms of Affordance
-Our survey reveals how users interact with ChatGPT and its feedback features, which influence the type and quality of data fed back into RLHF workflows.
-Respondents used ChatGPT frequently and in diverse ways.
-Over 71.6% (n=525) used ChatGPT at least weekly, and the most common use cases were answering specific questions (e.g., factual, technical, or theoretical questions), generating text (e.g., drafting emails, writing essays or reports), and academic research assistance (e.g., generating citations, summarising articles).
-This usage pattern is significant because a person cannot directly influence ML workflows outside that system.
+4.1.1. Request and Demand
+The ChatGPT interface requests feedback formatted in the “helpful, honest, and harmless” framework, as evident in the feedback submission options presented to users.
+This framing guides users toward specific types of evaluative feedback aligned with OpenAI’s stated development priorities.
+More fundamentally, the interface demands interaction with the system before allowing participation through one of the three modes for feedback submission.
 A user must prompt the dialogue agent before they can submit feedback.
-This design feature effectively demands interaction with an LLM before allowing participation through feedback submission, establishing the first key mechanism of affordance we observed.
-4.1.1. Submitting Feedback
+This design feature effectively makes system usage a prerequisite for participation in the feedback process, establishing a key mechanism of affordance we observed.
+Table 2. Types of Feedback Submitted (nn=526)
+Feedback Type Submitted Not Submitted
+Thumbs-up/down feedback 76.0% 24.0%
+Open-text feedback 29.0% 71.0%
+Feedback after regenerating response 62.9% 37.1%
+4.1.2. Encourage
+Our survey reveals several patterns in what forms of feedback the interface encourages.
+As shown in Table 2, a large majority of respondents (84.2%) submitted at least one form of feedback through the ChatGPT interface.
+Thumbs-up/down feedback was the most common, followed by feedback after regenerating responses, while open-text feedback was less common.
+Furthermore, among those who submitted feedback, respondents submitting thumbs-up/down feedback reported doing so most frequently (28.8% every/most times), followed by feedback after regenerating a response (27.8% every/most times), while respondents submitting open-text feedback reported doing so least frequently (15.2% every/most times) (see Table 3).
+The observed preference for simpler feedback features may be reinforced by the interface design itself, where the thumbs-up/down feedback icons are visually prominent after each response, while accessing the open-text feedback requires an additional click.
+Table 3. Frequency of Feedback Submission by Type
+Frequency Thumbs-up/down Open-text After Regenerating
+(nn=400) (nn=152) (nn=331)
+Every/Most times 28.8% 15.2% 27.8%
+About half the time 24.5% 11.2% 19.6%
+Sometimes 34.8% 50.7% 36.3%
+Rarely 12.0% 23.0% 16.3%
+Total 100.0% 100.0% 100.0%
+The interface appears to particularly encourage performance-related feedback.
+Respondents who submitted open-text feedback were primarily motivated by improving the performance and capabilities of the system.
+We asked this group of respondents to reflect on the most recent time they submitted open-text feedback and report their motivations for doing so.
+Respondents could select multiple motivations, but those who selected multiple motivations were forced to choose their primary motivation in the following question.
+Table 4 shows the primary motivations in order, taking the weighted frequencies.
+Table 4. Primary Motivations for Submitting Open-text Feedback (nn=152)
+Motivation Weighted Frequency
+To contribute to AI development 159
+To comment on use-case specific feedback 94
+To report a technical issue or bug 74
+To suggest new features or improvements 73
+To express social or ethical concerns about ChatGPT 32
+Other 15
+Respondents were also asked what feedback they thought most useful for improving ChatGPT.
+Based on our thematic analysis, respondents emphasised technical issues (bugs, glitches) and system performance (efficiency, speed).
+Concerns about accuracy (“hallucinations” or false information) and the reliability or truthfulness of responses were less common, while social and ethical concerns (harmful, biased, or misleading responses) were mentioned even less frequently.
+In summary, while the feedback features in the ChatGPT interface allow for open-text feedback, including commentary on social and ethical issues, our respondents predominantly focused on performance-related concerns when submitting feedback.
+This focus appears to be shaped by the affordances of the feedback system itself rather than by explicitly discouraging social or ethical feedback.
+Despite OpenAI’s initial blog post requesting feedback on “harmful outputs” and “novel risks”, our survey respondents appear to have interpreted the purpose of their participation primarily through the lens of improving ChatGPT’s performance.
+4.1.3. Discourage and Refuse
 The feedback features in the ChatGPT interface are designed to elicit individual responses, which effectively discourages collective feedback or discussions among users.
-This individualisation extends to the permanence of feedback: users cannot modify their input once submitted, even if they gain new information or their perspective changes.
+This individualisation extends to the permanence of feedback: the interface refuses modification of input once submitted, even if users gain new information or their perspective changes.
 While users could theoretically input the same prompt a second time and submit feedback, the stochastic nature of LLMs means each response is likely to be unique, making direct comparisons or revisions challenging.
 This design choice reinforces the individual nature of the feedback process and limits opportunities for evolving, collective input.
-Our survey findings indicate that the ChatGPT interface encourages thumbs-up/down feedback and feedback after regenerating a response.
-Most respondents (84.2%, n=526) indicated they had submitted feedback through one of the three methods in the ChatGPT interface.
-Thumbs-up/down feedback was the most common (76.0%, n=526), followed by feedback after regenerating responses (62.9%, n=526), while open-text feedback was less common (29.0%, n=526).
-Respondents submitting thumbs-up/down feedback also reported doing so most frequently (53.3% every, most, or about half the times they had used ChatGPT, n=400), followed by feedback after regenerating a response (47.4% every, most, or about half the times they had used ChatGPT, n=331).
-Respondents submitting open-text feedback reported doing so least frequently (26.4% every, most, or about half the times they had used ChatGPT, n=152).
-Respondents who submitted open-text feedback (n=152) were primarily motivated by improving the performance and capabilities of the system.
-We asked this group of respondents to reflect on the most recent time they submitted open-text feedback and report their motivations for doing so.
-Respondents could select multiple motivations, but those who selected multiple motivations were forced to choose their primary motivation in the following question.
-Taking the weighted frequencies, the primary motivations were (in order) to: “contribute to AI development”, “comment on use-case specific feedback”, “report a technical issue or bug”, and “suggest new features or improvements” (see Table 12 in Appendix A).
-Expressing social or ethical concerns about system outputs was the least common motivation (excluding the range of individual responses in the “other” option).
-4.1.2. Reflecting on Feedback
-We also asked respondents who submitted open-text feedback what feedback they thought was most useful for improving ChatGPT’s performance or impact.
-The responses revealed a hierarchy of feedback themes.
-The most prominent themes were technical issues (bugs, glitches, incorrect information) and system performance (efficiency, speed, and quality of responses).
-Less common but still significant were concerns about accuracy (‘hallucinations’ or false information) and the reliability and truthfulness of responses.
-Though less frequent again, some users highlighted social and ethical concerns (including harmful, biased, or misleading responses).
-Finally, some feedback focused on suggesting new features or improvements.
-Most respondents were ambivalent about whether their reasons for submitting feedback had been addressed—47.4% neither agreed nor disagreed, 23.7% somewhat agreed, and 16.4% somewhat disagreed (n=152).
-Few of those who submitted open-text feedback indicated they had received follow-up communication or updates from OpenAI (11.2%, n=152), though the vast majority of those who had indicated they were satisfied with the response or resolution.
-The feedback features in the ChatGPT interface allow for open-text feedback, including commentary on social and ethical issues.
-However, our survey results reveal that users predominantly focus on performance-related concerns when submitting feedback.
-This focus appears to be shaped by the affordances of the feedback system itself rather than by explicitly discouraging social or ethical feedback.
-Despite OpenAI’s initial blog post requesting feedback on risks and harms, users who responded to our survey seem to have interpreted the purpose of their participation primarily through the lens of improving ChatGPT’s performance.
-This pattern highlights how the design of feedback features can effectively constrain user perspectives and shape the nature of user input.
+Most respondents were ambivalent about whether their reasons for submitting feedback had been addressed—47.4% neither agreed nor disagreed, 23.7% somewhat agreed, and 16.4% somewhat disagreed (nn=152).
+Few of those who submitted open-text feedback indicated they had received follow-up communication or updates from OpenAI (11.2%, nn=152), though the vast majority of those who had indicated they were satisfied with the response or resolution.
+4.1.4. Allow
+While the ChatGPT interface may constrain direct feedback, it allows for alternative channels of expression.
+Most respondents (80.4%, nn=526) did not indicate they shared feedback about ChatGPT through other channels or platforms.
+However, those who did primarily used X/Twitter (42%), followed by Facebook (18%) and Reddit (17%).
+Several respondents engaged with dedicated communities, including subreddits like r/ChatGPT and online forums like the OpenAI Developer Forum.
+Most respondents who indicated they had shared feedback through other channels or platforms (71.9%, nn=103) reported they were likely to continue to share feedback through those channels.
 4.2. Conditions
+The mechanisms described above operate through three conditions: perception, dexterity, and cultural and institutional legitimacy.
+Each condition modifies how mechanisms function for different users, creating variation in how the interface affords participation.
 4.2.1. Perception
 Respondents generally perceived the feedback features in the interface as accessible and user-friendly.
-Over three-quarters reported confidence in their ability to evaluate and submit feedback about the performance and/or impact of generative AI systems such as ChatGPT (“Strongly Agree” (21.8%) or “Somewhat Agree” (56.5%), n=524).
-Of those who submitted at least one of the three feedback modes, almost all found doing so through the features extremely or somewhat easy (95.1% n=442).
+As shown in Table 5, over three-quarters reported confidence in their ability to evaluate and submit feedback about the performance and/or impact of generative AI systems such as ChatGPT.
+Of those who submitted at least one of the three feedback modes, almost all found doing so through the features extremely or somewhat easy (95.1% nn=442).
+Table 5. Confidence Submitting Feedback (nn=524)
+Confidence level Percentage
+Strongly agree 21.8%
+Somewhat agree 56.5%
+Neither agree nor disagree 15.1%
+Somewhat disagree 5.7%
+Strongly disagree 1.0%
+Total 100.0%
 However, some users were more confident than others in submitting feedback.
-Younger users (18-34) and those who frequently engaged with AI-related topics expressed more confidence than would be expected if there was no relationship between these factors and confidence.
+Those who frequently engaged with AI-related topics expressed more confidence than would be expected if there was no relationship between these factors and confidence (weak association; see Table 7 for all statistically significant relationships).
 Conversely, users with less frequent engagement were less confident than would be expected.
-We found weak associations between confidence and age ($\chi^{2}$ = 17.068, adj.
-$p$-value = 0.045, df = 6, Cramér’s V = 0.131) and frequent engagement with AI-related topics ($\chi^{2}$2 = 20.726, adj.
-$p$-value = 0.045, df = 8, Cramér’s V = 0.144).
+Younger users (18-34) also showed a weak association with higher confidence in the initial analysis, though this finding was sensitive to the statistical approach used; the association lost significance when tested against the original five-point scale.
+The differences in perception impact which mechanisms operate for different user groups.
+For confident users familiar with AI topics, the interface appears to encourage feedback submission through its accessible design.
+For the 21.8% of respondents who did not express confidence in evaluating and submitting feedback (those responding “Neither Agree nor Disagree” to “Strongly Disagree”), these same interface features may effectively discourage participation despite their availability.
 4.2.2. Dexterity
-Few respondents reported barriers to accessing ChatGPT or submitting feedback, suggesting that users with varying levels of dexterity can provide input.
-When asked about barriers to accessing ChatGPT or submitting feedback, most respondents (77.3%, n=519) reported experiencing no barriers.
+Most respondents reported no barriers to accessing ChatGPT or submitting feedback (77.3%, nn=519), suggesting that users with varying levels of dexterity can provide input.
 Among those who encountered barriers, the most common were language and internet connectivity issues.
 We did not find significant associations between most demographic factors or AI experience and knowledge factors and those who experienced barriers.
-While those working in AI were more likely to report experiencing barriers than would be expected if there was no relationship between the variables, the association was weak ($\chi^{2}$ = 9.821, adj.
-$p$-value = 0.023, df = 1, Cramér’s V = 0.143).
-However, the requirement to prompt the dialogue agent before submitting feedback means only users can submit feedback through the interface’s features.
+While those working in AI were more likely to report experiencing barriers than would be expected if there was no relationship between the variables, the association was weak (see Table 7).
+However, as noted earlier, the requirement to prompt the dialogue agent before submitting feedback effectively demands system usage as a prerequisite for participation and refuses feedback from non-users.
 Our sample, therefore, does not include non-user perspectives, and given that most respondents used ChatGPT at least weekly, it also includes limited input from infrequent users.
-Notably, of those who completed the survey but did not indicate they had submitted feedback through one of the three options, the most common reason was that they were unaware of the feedback option (34.6%, n=78).
-While the ChatGPT interface may constrain direct feedback, respondents found alternative ways to express their opinions and participate in discussions about the system.
-Most respondents (80.4%, n=526) did not indicate they shared feedback about ChatGPT through other channels or platforms.
-However, those who did primarily used X/Twitter (42%), followed by Facebook (18%) and Reddit (17%).
-Several respondents engaged with dedicated communities, including subreddits like r/ChatGPT and online forums like the OpenAI Developer Forum.
-Most respondents who did indicate they had shared feedback through other channels or platforms (71.9%, n=103) reported they were likely to continue to share feedback through those channels.
+Notably, of those who completed the survey but did not indicate they had submitted feedback through one of the three options, the most common reason was that they were unaware of the feedback options (see Table 6).
+For these users, what designers intended to allow is functionally refused.
+Table 6. Main Reasons for Not Submitting Feedback (nn=78)
+Reason Count Percentage
+Unaware of feedback option 27 34.6%
+Not enough time or interest in submitting feedback 21 26.9%
+Satisfied with performance and have not felt need to submit feedback 17 21.8%
+Unsure how to provide feedback 7 9.0%
+Other 5 6.4%
+Unsatisfied with performance, but do not believe feedback will lead to improvements 1 1.3%
+Total 78 100.0%
+Also, as noted earlier, some respondents found alternative channels and platforms (outside the ChatGPT interface) to express their opinions and participate in discussions about the system.
+These results suggest that while the interface discourages certain forms of participation (e.g., collective feedback or discussions among users), users with sufficient dexterity sought platforms that allow forms of engagement the original interface refuses.
 4.2.3. Cultural and Institutional Legitimacy
-Interactive feedback features may appeal more to users with certain cultural and institutional backgrounds, which can influence who submits feedback and the types of feedback they provide.
-Indeed, most of our survey respondents were young and identified as men, which reflects ChatGPT’s reported user base (Vogels, 2023) and other studies of contributors to human feedback datasets (Kirk et al., 2024).
-Generally, feedback types were proportional across demographic groups, though we found that respondents’ region had a moderate association with providing open-text feedback ($\chi^{2}$ = 13.990, adj.
-$p$-value = 0.023, df = 3, Cramér’s V = 0.192).
-Instead, our survey results suggest that feedback features appeal more to users with frequent engagement or familiarity with AI, ML, and data science.
-Frequent ChatGPT users, particularly daily users, were more likely to provide thumbs-up/down feedback ($\chi^{2}$ = 13.689, adj.
-$p$-value = 0.045, df = 4, Cramér’s V = 0.166) than would be expected if there was no relationship between the variables, and those who previously used AI tools ($\chi^{2}$ = 8.926, adj.
-$p$-value = 0.023 df = 1, Cramér’s V = 0.154) were more likely to provide open-text feedback than would be expected.
-We also observed moderate associations suggesting more frequent ChatGPT users and those often engaging with AI-related topics were more likely to submit open-text feedback than would be expected if there was no relationship between these variables.
-However, these associations did not reach statistical significance after adjusting $p$-values.
-Additionally, we noted a borderline significant association between education or training in AI and open-text feedback submission (adj.
-$p$-value = 0.059).
-4.3. Summary of Affordances
+Feedback features may appeal more to users with certain cultural and institutional backgrounds, which can influence who submits feedback and the types of feedback they provide.
+Indeed, most of our survey respondents were young and identified as men, consistent with ChatGPT’s reported user base (Vogels, 2023) and studies of contributors to human feedback datasets (Kirk et al., 2024).
+Generally, feedback types were proportional across demographic groups, though geographic region moderately correlated with open-text feedback submission (see Table 7).
+More significantly, the survey insights in Table 7 suggest that feedback features appeal more to users with frequent engagement or familiarity with AI, ML, and data science.
+Frequent ChatGPT users, particularly daily users, were more likely to provide thumbs-up/down feedback and those who previously used AI tools were more likely to provide open-text feedback than would be expected if there was no relationship between the variables.
+We also observed that education in AI correlated with open-text feedback submission, though this relationship fell just short of statistical significance (adj.
+$p$-value = 0.059, see Table 38).
+These patterns suggest that technical background, rather than demographic characteristics, most strongly influenced how our respondents engaged with feedback features.
+These insights suggest how cultural and institutional legitimacy conditions shape the operation of affordance mechanisms.
+While the interface allows feedback from any user, it appears to more successfully encourage feedback from those whose technical knowledge or experience aligns with existing institutional norms in AI development.
+Table 7. Significant Associations Between User Characteristics and Feedback Behaviours
+Relationship $\chi^{2}$ $p$-value Cramér’s V Interpretation
+Thumbs-Up/down and After Regenerating 50.419 ¡.001 .319 Moderate
+Education/Training in AI and Confidence 12.773 .023 .160 Weak
+Working in AI and Barriers 9.821 .023 .143 Weak
+Region and Open-text feedback 13.990 .023 .192 Moderate
+Prior AI use and Open-text feedback 8.926 .023 .154 Weak
+Thumbs-Up/down and ChatGPT frequency 13.689 .045 .166 Weak
+AI topics frequency and Confidence 20.726 .045 .144 Weak
+Age and Confidence 17.068 .045 .131 Weak
+Table 8. Summary of Affordances
+Feature(s) Mechanism of Affordance Outcomes For Whom/Condition
+Checkboxes and structure in feedback windows Request Feedback formatted in the “helpful, honest, and harmless” framework Feedback from users to developers (Dexterity)
+Login and interaction requirement Demand System usage as prerequisite for participation Excludes non-users (Dexterity)
+Thumbs-up/down buttons; Regeneration feedback options; Open-text box with categorical selection options Encourage Simple, frequent feedback; Performance-focused feedback; Technical issue reporting For users with AI/ML experience and knowledge (Cultural and institutional legitimacy); For frequent users familiar with the system (Perception, Dexterity)
+Single-user login Discourage Collective feedback; Discussion among users Excludes collectives of users or stakeholders (Cultural and institutional legitimacy)
+No edit functionality after submission; Stochasticity of LLMs Refuse Modification of feedback over time; Comparative evaluation of responses For all users (Dexterity)
+Access to external platforms Allow Alternative channels for feedback and discussion For technically knowledgeable users (Dexterity); For users with social media presence (Cultural and institutional legitimacy)
+4.3. Synthesis: How Affordances Shape Participation
 Table 8 outlines the affordances of the feedback features in the ChatGPT interface based on the analysis above, organised by the six mechanisms proposed by the framework.
-The affordances are linked to the conditions (perception, dexterity, and cultural and institutional legitimacy) that shape how users interact with and perceive the feedback features.
-Table 1. Summary of Affordances
-Mechanism Affordance Condition
-Request Feedback formatted in the ‘helpful, honest, and harmless’ framework Feedback from users to developers (Dexterity)
-Demand Interaction with ChatGPT before feedback submission Status as a user (Dexterity)
-Encourage Thumbs-up/down feedback; Feedback after regenerating a response; Performance-related open-text feedback Ease of use and accessibility of interface (Perception); Frequency of use and familiarity with the system (Perception, Dexterity); AI/ML experience and knowledge (Cultural and institutional legitimacy)
-Discourage Feedback from collectives; Discussions among users Individualised feedback design (Cultural and institutional legitimacy)
-Refuse Modifying feedback over time Stochastic nature of the system (Dexterity)
-Allow Alternative channels for feedback and discussion (e.g., social media, online forums) Availability of alternative channels (Dexterity); Accessibility and appeal of feedback features to certain user groups (Cultural and institutional legitimacy)
+The affordances are linked to the conditions that shape how users interact with and perceive the feedback features.
+In summary, the interface requests feedback formatted in the “helpful, honest, and harmless” framework, demands interaction before feedback submission, and encourages simple, frequent feedback focused on performance improvement.
+However, the interface discourages collective feedback and discussions among users, while allowing alternative channels for such discussions, and refuses feedback modification over time.
+Additionally, our survey indicates that the feedback features may appeal more to users with frequent engagement, prior experience, and knowledge of AI, ML or data science.
 5. Discussion
 Participation is never fully open or anarchic; it is always formatted with “specific rules and scripts” (Kelty, 2020).
-Feedback features in interfaces to LLMs format user input into RLHF workflows.
-Therefore, the design of these features shapes the legitimacy of user participation in the iteration of LLMs.
-Our analysis of feedback features in ChatGPT suggests that the feedback format imposed by the interface is simple, individualised, unidirectional, and focused on performance-related concerns.
-This format significantly constrains user participation and risks becoming self-perpetuating—users may internalise limited roles in the iteration of LLMs, discouraging them from seeking broader engagement or viewing themselves as co-creators in LLM development.
+Feedback features in interfaces to LLMs format user input into RLHF workflows, which both legitimises user participation in the iteration of LLMs and constrains such participation to unidirectional channelling of atomised, de-contextualised user preferences.
+The interface design also channels user input toward easily processed feedback focused on performance-related concerns; 76% of our respondents used thumbs-up/down feedback, but only 29% provided open-text feedback, with technical or performance motivations significantly outweighing social or ethical concerns.
+This highlights a clear tension between technology companies’ stated intentions (e.g., OpenAI initially sought diverse, substantive feedback to uncover potential risks and harms (OpenAI, 2022)) and how these features are actually used (simple, performance-focused feedback).
+While ostensibly democratising participation, we ask whether participation through feedback features is in name only, part of a “proceduralized participation… [that] can often be implemented without any challenge to existing power relations?” (Kelty, 2020).
+Our sample primarily comprises young, tech-savvy early adopters of ChatGPT.
+While this limits generalisability, the demographic skew itself highlights how affordances of feedback features in initial deployments of LLMs might privilege certain groups of users, and the narrow forms of participation we observed among these users demonstrate how the interface’s affordances encouraged simple, performance-focused feedback in this context.
+These constraints among early users raise broader questions about how such feedback processes might function as systems scale.
+If early adopters primarily provide simple, performance-focused feedback due to the interface design, users with less technical experience and knowledge, or facing other barriers, might find it even more difficult to articulate complex social or ethical concerns through these channels.
+This indicates the potential for a self-perpetuating dynamic where users internalise limited roles as participants in LLM iteration.
 This phenomenon is not unique to LLMs; similar patterns have been observed in other technological domains, such as social mediaplatforms (Gillespie, 2020).
 Affordance theory as a design tool requires an orientation that “technologies, and the worlds they build, need not be as they are and can, instead, be otherwise” (Davis, 2023).
-As LLMs increasingly become infrastructure, it is important to consider participation beyond mere preference elicitation.
-In this section, we discuss two directions for the (re)design of feedback processes.
-Initially, we review recent developments in deliberative approaches to feedback, such as recent democratic experiments of OpenAI and Anthropic.
-However, we argue that users and stakeholders can never get to the heart of closed systems by commenting on model outputs.
-Instead, we encourage a shift in approach from alignment of model outputs to specific user preferences to dialogue with the public about the purpose and applications of LLMs, and attention to the ongoing infrastructuring work required to support and sustain diverse ‘publics’ over time.
+Drawing on affordance theory, we discuss two potential directions for the (re)design of feedback processes to move beyond the constraints revealed by our analysis, aiming at shifting participation from procedural to more substantive forms.
 5.1. Deliberative Processes
-As previously noted, the existing feedback features request feedback from users to developers and discourage both feedback from groups of users and discussion amongst users.
-In other words, feedback through existing features is atomised and unidirectional (Wang et al., 2023).
-Additionally, as the number of ChatGPT users has grown and the system has stabilised over time, user input has been constrained from a wide to a narrow aperture—from free-text feedback forms to open-text feedback windows to singular buttons focused on crowdsourcing inputs to RLHF.
-These latest feedback formats are inadequate for resolving social and ethical considerations, requiring us to consider other design approaches.
-If algorithmic fairness is primarily a political task, as Wong (Wong, 2020) argues, rather than solely a technical task, we must consider how to resolve issues of bias and fairness politically rather than technically.
+As discussed in Section 2.1, resolving social and ethical issues in AI deployment is increasingly framed as a political task rather than a technical task.
+One potential direction for redesigning feedback processes to address such issues politically is to incorporate elements of deliberation—structured dialogue among users and (potentially) with developers.
 Alongside aggregating citizen preferences through voting, deliberation among citizens is a key element of democratic legitimacy (Fishkin, 1991; Knight and Johnson, 1994; Curato et al., 2017).
-Deliberative democracy involves an association of members who govern their affairs by deliberation among the members (Elster and Przeworski, 1998)—making collective decisions determined by those subject to the decisions: not only their preferences but also their reasoning (Cohen, 2005).
-Such associations tend to be between broad groups of people—in this context, users, developers, and potentially groups of both—and are directed towards participants considering others’ perspectives and potentially changing their own over time.
-Interestingly, survey respondents identified several other platforms on which they discuss their experiences and preferences regarding ChatGPT, engaging in the kind of collective deliberation that the existing feedback features discourage.
-For example, OpenAI discussion forums allow communication between users and developers and subreddits such as r/ChatGPT allow discussion amongst users (without monitoring or oversight by OpenAI).
-However, it is unclear whether or how these discussions are input into development processes; while they function as evaluative feedback, there is no clear cybernetic feedback loop.
+Deliberative democracy involves an association of members who govern their affairs through deliberation among the members (Elster and Przeworski, 1998)—making collective decisions based not only on their preferences but also on their reasoning (Cohen, 2005).
+Such associations typically involve broad groups of people and encourage participants to consider others’ perspectives and potentially revise their own views over time
+Deliberative processes could change the existing affordances of user participation in iterating LLMs.
+For example, where current interfaces discourage reasoned discussion among users, deliberative platforms could explicitly request reasoning from participants and allow for bidirectional communication between users and developers, along with evolving perspectives, moving toward genuine feedback loops rather than the unidirectional channels currently prevalent.
+Interestingly, some respondents identified several other platforms on which they discuss their experiences and preferences regarding ChatGPT, engaging in the kind of collective deliberation that the existing feedback features discourage.
+For example, OpenAI discussion forums allow communication between users and developers and subreddits such as r/ChatGPT allow discussion among users (without monitoring or oversight by OpenAI).
+However, it is unclear whether or how these discussions feed back into development.
 Some technology companies have been experimenting with deliberative modes for feedback about AI systems.
 Recent efforts by OpenAI to fund research into technical solutions for “democratic inputs into AI” (Eloundou and Lee, 2024) and experiments by Anthropic to draft a constitution for Claude through deliberative processes (Huang et al., 2024) have attracted significant attention.
-These efforts go beyond eliciting and aggregating user preferences towards encouraging deliberation amongst people with diverse perspectives (including non-users) and demonstrating how the preferences and reasoning of multiple participants are input (as cybernetic feedback) into development workflows.
+These efforts go beyond eliciting and aggregating user preferences.
+They encourage deliberation among users and (potentially) other stakeholders, while demonstrating how multiple participants’ preferences and reasoning feed into development workflows.
 In addition, Meta has experimented with Community Forums to deliberate on the rulesets for Generative AI (Broxmeyer, 2024), and Google has developed the STELA process for deliberative discussions with historically underrepresented groups in the U.S.A.
@@
 While these processes address some of the limitations of existing feedback features, they remain largely experimental and peripheral to the dominant method of feedback submission through browser-based interfaces.
-Most of these approaches (with the Google example as an exception) take as a given that consensus on the output of the process is desirable and achievable.
-However, refusal, agonism and dissensus are key elements of democratic processes (DiSalvo, 2015, 2010; Zong and Matias, 2024).
-When consulting multiple and diverse communities, values and preferences for AI design may be in conflict (Sambasivan et al., 2021; Qadri et al., 2023) and difficult to reconcile (Santurkar et al., 2023; Durmus et al., 2023; Miller et al., 2007).
-This is particularly the case where the processes are managed by those with commercial interests in consensus outcomes, reflecting long-standing concerns in participatory design (Sanders and Stappers, 2014).
-Moreover, such democratic inputs to the alignment problem, much like the existing feedback processes they are intended to augment or replace, narrowly scope participation around the output of specific families of models rather than engaging participants in earlier stages of the development process.
-Although deliberative processes may address some limitations of existing mechanisms and change how companies like OpenAI and Anthropic make decisions about model behaviour, they do not significantly alter the power dynamics between users, the public, and platforms.
-Users and other participants remain as consumers rather than co-creators, limited in their influence over core aspects of earlier stages of development.
+Most of these approaches (with the STELA process as an exception) take as a given that consensus on the output of the process is desirable and achievable.
+However, refusal, agonism and dissensus are all key elements of democratic processes (DiSalvo, 2015, 2010; Zong and Matias, 2024).
+When consulting multiple and diverse communities, values and preferences for AI design may be in conflict (Sambasivan et al., 2021; Qadri et al., 2023) and difficult to reconcile (Santurkar et al., 2023; Durmus et al., 2024; Miller et al., 2007).
+Such difficulty is particularly acute where the processes are managed by those with commercial interests in consensus outcomes, reflecting long-standing concerns in participatory design (Sanders and Stappers, 2014).
+Moreover, such democratic inputs to the alignment problem, much like the existing feedback features they are intended to augment or replace, narrowly scope participation around the output of specific families of models, refusing participants meaningful engagement with earlier development stages (Delgado et al., 2023).
+Although potentially addressing some limitations of current feedback features, users and stakeholders remain as consumers rather than co-creators.
 5.2. Infrastructuring Participation
-Our study reveals a tension between technology companies’ stated intentions and their implementation of feedback features in LLMs.
-ChatGPT exemplifies this disparity: what OpenAI initially sought (diverse, substantive feedback to uncover potential risks and harms) gave way to what they could effectively use within their existing workflows (easily processed, formatted data).
-This change not only constrains user input but also, as our survey indicates, favours frequent ChatGPT users and those with AI experience, narrowing the group of users influencing LLM iteration.
-While eliciting user preferences as input into RLHF workflows may be well-intentioned and useful for technology companies to improve model performance, it is inadequate for the larger task facing us all.
-We all face the challenge of publicly, deliberatively, and inclusively addressing the ongoing tectonic shifts in information and knowledge generation across economic, social, cultural, and political domains before LLMs become unremarkable and unquestionable infrastructure.
-If we are to shift our focus from aligning model outputs to specific user preferences towards dialogue with the public about the purpose and applications of LLMs, then we must think critically about who will form this ‘public’, how it will form, and what channels facilitate its participation.
-Currently, technology companies configure participation from atomised users of dialogue agents like ChatGPT, tightly defining processes for feedback on model outputs.
-In ChatGPT’s case, the ‘public’ is constructed primarily to address a matter of concern to OpenAI rather than those engaged as participants.
-Moreover, groups or institutions affected by the development and deployment of LLMs lack direct mechanisms to influence these feedback processes (Wang et al., 2023).
-Technology companies might argue for increasing the number and diversity of participants providing feedback through existing channels.
-However, this approach misses the core issue.
-Drawing parallels with Arnstein’s (Arnstein, 1969) critique of participation in the context of urban planning, we argue that identifying problems in the proceduralisation of participation through feedback features in LLMs is not a call for more involvement but for a fundamental rethinking of the process.
-We echo Wang et al.
-(Wang et al., 2023) in urging companies to focus on the participation of groups or institutions representing populations likely to be significantly impacted by a given LLM.
-Infrastructuring such participation requires deep engagement with the methodological tools of Participatory Design (PD), which has a long-standing practice of critical thinking about how to form and sustain ‘publics’ to address participants’ concerns over time (DiSalvo, 2009; Le Dantec and DiSalvo, 2013).
-When implemented effectively, PD prioritises future issues concerning participants rather than solely focusing on the immediate concerns of designers or researchers.
-Recent research offers promising directions for rethinking processes and infrastructure for participation in LLM development.
-The STELA process (Bergman et al., 2024) demonstrates one method for conducting community-centred deliberative discussions about LLMs, albeit focused on aligning models with the preferences of specific groups of people.
+While our insights are based on early adopters, the constraints revealed by our analysis highlight the limitations of current processes for feedback on harmful model outputs or risks from LLM deployment.
+Existing processes appear to be efficient channels for atomised preferences from specific user groups, but insufficient conduits for dialogue on social and ethical issues regarding LLMs.
+Echoing Arnstein’s (Arnstein, 1969) critique of participation in the context of urban planning, identifying problems in the proceduralisation of participation through existing processes is not simply a call for more involvement, but instead requires a fundamental rethinking of the process.
+We advocate for moves beyond aligning model outputs to specific user preferences toward feedback processes that embrace productive disagreement and plurality.
+As DiSalvo (DiSalvo, 2010) argues, designers can enhance democratic engagement by “creating and enabling … spaces of contest” where “difference and dissensus are brought forward and the assumptions and actions that shape power relations and influence are revealed and challenged” (DiSalvo, 2010).
+To create feedback channels capable of supporting these more contested forms of engagement requires actively “infrastructuring” participation—creating and sustaining social, technical, and institutional structures with different affordances to address matters of concern to participants (Le Dantec and DiSalvo, 2013).
+Such infrastructure must aim at changing the conditions of participation, to actively encourage engagement from stakeholders affected by LLM deployments, including non-users and those without technical backgrounds, and allow forms of input that address participants’ concerns beyond immediate systemperformance.
+Recent research offers promising practical directions.
+The STELA process (Bergman et al., 2024) demonstrates one method for conducting community-centred deliberative discussions about LLMs with marginalised groups who may or may not be users of LLMs, albeit focused on aligning models with the preferences of those groups of people.
 Suresh et al.
-(Suresh et al., 2024) expand on this concept by proposing a comprehensive framework for domain-specific community infrastructure, including shared technical resources, norms, and governance structures.
-Their approach aims to enable more meaningful participation by groups and institutions in shaping the development and application of foundation models in specific domains.
-Building on these ideas, we encourage participatory workshops with stakeholders in fields like education, healthcare, and journalism as a practical mechanism for realising and refining such domain-specific infrastructure.
-However, implementing these participatory approaches faces significant challenges: maintaining local relevance while extending across domains and regions, translating insights into actionable development inputs, and securing sustainable funding for ongoing infrastructure in the rapidly changing landscape of AI development.
-To advance meaningful public participation in the iteration of LLMs, the focus of research and effort must be on addressing these challenges rather than defaulting to procedural and perfunctory engagement.
+(2024) expand on this concept by proposing a comprehensive framework for domain-specific community infrastructure, including shared technical resources, norms, and governance structures.
+Their approach aims at enabling participation by stakeholder groups and institutions to shape the development and application of foundation models in specific domains.
+Building on these ideas, we encourage participatory design workshops with stakeholders, in fields like education, healthcare, and journalism, that use techniques encouraging agonism (see, e.g., (Björgvinsson et al., 2012)) to surface conflicts and refine such domain-specific infrastructure.
+Building and sustaining this infrastructure is not a one-off task; it requires ongoing effort to secure resources, build trust with participants, navigate conflict, ensure local relevance while extending across domains and regions, and translate outputs into actionable development inputs (Le Dantec and DiSalvo, 2013).
+However, these are not reasons to default to simpler, constrained modes of participation.
+If the goal of feedback processes is to afford meaningful public influence over technologies with profound societal impact, then the focus of research and effort must be on addressing such challenges to infrastructuring sustained, pluralistic, and sometimes contentious dialogue with affected stakeholders.
 6. Conclusion
-Our analysis reveals that existing feedback features in the ChatGPT interface constrain participation around developers’ instrumental needs.
-This approach to user input and participation views engagement as a means to an end rather than an opportunity for dialogue.
-While deliberative approaches may address some limitations of existing mechanisms, the scope of user and non-user input remains constrained.
-To move beyond perfunctory engagement, we call for a shift in focus from user preference elicitation to building robust infrastructures for sustained, meaningful dialogue about the purpose and application of LLMs.
+Our analysis of ChatGPT’s feedback features reveals significant constraints on user participation, based on a survey of early adopters’ experiences.
+The interface appears to encourage simple, frequent, and performance-focused feedback while discouraging collective input and discussions among users.
+We argue that this constrained approach to participation suggests limitations for meaningful dialogue about social and ethical issues regarding LLM deployment as system usage becomes more widespread.
+Three research directions could substantiate and extend our analysis.
+Firstly, empirical studies could test whether deliberative processes better capture social and ethical concerns than current processes.
+Additionally, longitudinal research could track both how feedback patterns evolve as broader populations adopt LLMs, and how organisations frame their feedback processes over time.
+Such work might reveal growing divergences between companies’ stated commitments and actual practices.
+Most importantly, we encourage ethnographic and action research examining the practical work of infrastructuring.
+Ethnographic studies could document ongoing attempts to create participatory structures in specific domains, while action researchers could collaborate with affected communities to prototype and sustain new forms of engagement (see, e.g., (Tseng et al., 2025)).
+Without fundamental changes, the scope of user (and non-user) input will remain constrained.
+To move beyond perfunctory engagement requires shifting focus from user preference elicitation to building participatory infrastructures with affordances that allow and encourage sustained dialogue—including disagreement—with affected stakeholders about the purpose and applications of LLMs.
 As these technologies become increasingly embedded in information, workplace, and entertainment services, the urgency of this task cannot be overstated.
-Appendix A Survey Results
-Appendix A compiles the survey results from our study.
-It contains a series of tables with detailed breakdowns of respondent usage patterns, feedback behaviours, demographics, and perceptions of submitting feedback using the interactive features in the ChatGPT interface.
+Acknowledgements.
+This research is supported by an Australian Government Research Training Program (RTP) Scholarship and a Florence Violet McKenzie Scholarship.
+We thank the organisers and participants of the CHI 2023 Workshop on Designing Technology and Policy Simultaneously and the 2023 AusSTS Conference for feedback on earlier drafts.
+We also thank Professor Alice Richardson of the Statistical Support Network, ANU, for helpful advice on the statistical analysis in this article.
+Finally, we thank our reviewers for their feedback and suggestions.
+Appendices
+Appendix A Survey Questions
+Appendix A lists the substantive questions from the online survey instrument used for this study.
+Screening
+  1. (1)
+    Have you used ChatGPT in any capacity? e.g., to generate text, answer a question, or simply experiment with the system.
+    • Yes
+    • No
+ChatGPT Usage
+  1. (2)
+    How often have you used ChatGPT since its release in November 2022?
+    • Daily
+    • A few times a week
+    • Once a week
+    • Monthly
+    • Rarely
+  2. (3)
+    In which of the following ways have you used ChatGPT? Please select all that apply.
+    • Academic research assistance (e.g., generating citations, summarising articles)
+    • Generating text (e.g., drafting e-mails, writing essays or reports)
+    • Answering specific questions (e.g., factual, technical, or theoretical questions)
+    • Creative writing (e.g., stories, poetry, or brainstorming ideas)
+    • Language learning or translation
+    • Personal advice or emotional support (e.g., relationship or career advice)
+    • Professional or technical problem-solving (e.g., programming, troubleshooting)
+    • Entertainment or casual conversation (e.g., jokes, trivia, or random facts)
+    • Other (Specify)
+  3. (4)
+    In which ways have you used ChatGPT most frequently? Please order from 1 = most frequent, 2 = next most frequent, and so on.
+    • Academic research assistance (e.g., generating citations, summarising articles)
+    • Generating text (e.g., drafting e-mails, writing essays or reports)
+    • Answering specific questions (e.g., factual, technical, or theoretical questions)
+    • Creative writing (e.g., stories, poetry, or brainstorming ideas)
+    • Language learning or translation
+    • Personal advice or emotional support (e.g., relationship or career advice)
+    • Professional or technical problem-solving (e.g., programming, troubleshooting)
+    • Entertainment or casual conversation (e.g., jokes, trivia, or random facts)
+    • Other (Specify)
+Feedback Submission: Thumbs
+  1. (5)
+    Have you submitted feedback on ChatGPT’s responses using the up and down thumbs?
+    • Yes
+    • No
+  2. (6)
+    (If Yes to previous) How often have you submitted feedback using the up and down thumbs?
+    • Every time I have used ChatGPT
+    • Most times I have used ChatGPT
+    • About half the time I have used ChatGPT
+    • Some times I have used ChatGPT
+    • Rarely
+Feedback Submission: Open-Text
+  1. (7)
+    Have you submitted open-text feedback on ChatGPT’s responses?
+    • Yes
+    • No
+  2. (8)
+    (If Yes to previous) How often have you submitted open-text feedback?
+    • Every time I have used ChatGPT
+    • Most times I have used ChatGPT
+    • About half the time I have used ChatGPT
+    • Some times I have used ChatGPT
+    • Rarely
+Feedback Submission: Regenerate Response
+  1. (9)
+    Have you submitted feedback after regenerating one of ChatGPT’s responses?
+    • Yes
+    • No
+  2. (10)
+    (If Yes to previous) How often have you submitted feedback after regenerating a response?
+    • Every time I have regenerated a response
+    • Most times I have regenerated a response
+    • About half the time I have regenerated a response
+    • Some times I have regenerated a response
+    • Rarely
+Reasons for Not Submitting Feedback
+  1. (11)
+    (If never submitted any feedback) What is the main reason you have not submitted feedback through the ChatGPT interface?
+Please select the main reason only.
+NB: ‘Performance’ refers to the helpfulness, truthfulness, or harmlessness of responses from ChatGPT to a prompt.
+    • Unaware of feedback option
+    • Satisfied with the performance and have not felt the need to submit feedback
+    • Unsatisfied with the performance, but do not believe feedback will lead to improvements
+    • Unsure how to provide feedback
+    • Not enough time or interest in submitting feedback
+    • Other (Specify)
+Motivations
+  1. (12)
+    (If submitted open-text feedback) Think back to the most recent time you submitted feedback through the interface.
+What motivated you to submit feedback about ChatGPT?
+Please select all that apply.
+    • To report a technical issue or bug
+    • To suggest new features or improvements
+    • To express social or ethical concerns about ChatGPT
+    • To comment on use-case specific feedback
+    • To contribute to AI development
+    • Other (Specify)
+  2. (13)
+    (If selected multiple motivations previously) What was your main motivation for submitting feedback? Please place a 1 in the box next to your main motivation.
+    • To report a technical issue or bug
+    • To suggest new features or improvements
+    • To express social or ethical concerns about ChatGPT
+    • To comment on use-case specific feedback
+    • To contribute to AI development
+    • Other (Specify)
+  3. (14)
+    (If submitted open-text feedback) What kind of feedback do you think is most useful for improving ChatGPT’s performance or impact?
+e.g., reporting technical issues, suggesting new features, expressing social or ethical concerns, and so on.
+“Performance” refers to the helpfulness, truthfulness, or harmlessness of responses from ChatGPT to a prompt.
+“Impact” refers to changes in society and culture from ChatGPT.
+[Open Text]
+  4. (15)
+    (If submitted open-text feedback) Have you received any follow-up communication or updates regarding your feedback?
+    • Yes
+    • No
+  5. (16)
+    (If Yes to previous) How satisfied are you with the response or resolution?
+    • Extremely satisfied
+    • Somewhat satisfied
+    • Neither satisfied nor dissatisfied
+    • Somewhat dissatisfied
+    • Extremely dissatisfied
+  6. (17)
+    (If submitted open-text feedback) How much do you agree with this statement: “My reasons for submitting feedback about ChatGPT through the interface have been addressed by OpenAI.”
+    • Strongly agree
+    • Somewhat agree
+    • Neither agree nor disagree
+    • Somewhat disagree
+    • Strongly disagree
+Reflections
+  1. (18)
+    (If submitted any feedback) Was submitting feedback through the ChatGPT interface easy or difficult?
+    • Extremely easy
+    • Somewhat easy
+    • Neither easy nor difficult
+    • Somewhat difficult
+    • Extremely difficult
+  2. (19)
+    How likely are you to submit feedback through the ChatGPT interface in the future?
+    • Extremely likely
+    • Somewhat likely
+    • Neither likely nor unlikely
+    • Somewhat unlikely
+    • Extremely unlikely
+  3. (20)
+    (Conditional logic based on Q19) Why are you likely to submit feedback through the ChatGPT interface in the future? [Open Text]
+  4. (21)
+    (Conditional logic based on Q19) Why are you not likely to submit feedback through the ChatGPT interface in the future? [Open Text]
+Feedback via Other Channels
+  1. (22)
+    Have you provided feedback on ChatGPT’s performance or impact through other channels or platforms?
+e.g., social media platforms, online discussion forums, contacting OpenAI, and so on.
+“Performance” refers to the helpfulness, truthfulness, or harmlessness of responses from ChatGPT to a prompt.
+“Impact” refers to changes in society and culture from ChatGPT.
+    • Yes
+    • No
+  2. (23)
+    (If Yes to previous) Which other channels or platforms have you used to provide feedback about ChatGPT’s performance or impact? Please select all that apply.
+    • X (Please list any specific # or communities)
+    • Mastodon (Please list any specific # or servers)
+    • Linkedin (Please list any groups)
+    • Facebook (Please list any groups)
+    • Reddit forums (Please specify which subreddits e.g., r/ChatGPT)
+    • Discord servers (Please specify which servers)
+    • Other online discussion forums (Please specify e.g., Github Community, OpenAI Community)
+    • Contacting OpenAI (e.g., via e-mail)
+    • Other (Specify)
+  3. (24)
+    (If Yes to using other channels) How likely are you to use those channels or platforms to provide feedback on ChatGPT in the future?
+    • Extremely likely
+    • Somewhat likely
+    • Neither likely nor unlikely
+    • Somewhat unlikely
+    • Extremely unlikely
+  4. (25)
+    (Conditional logic based on Q24) Why are you likely to provide feedback on ChatGPT through those channels or platforms in the future? [Open Text]
+  5. (26)
+    (Conditional logic based on Q24) Why are you not likely to provide feedback on ChatGPT through those channels or platforms in the future? [Open Text]
+Demographics
+  1. (27)
+    What is your age?
+    • Under 18
+    • 18-24
+    • 25-34
+    • 35-44
+    • 45-54
+    • 55-64
+    • 65+
+    • Prefer not to say
+  2. (28)
+    How would you describe your gender?
+    • Man
+    • Woman
+    • Non-binary/third gender
+    • Prefer to self-describe (Specify)
+    • Prefer not to say
+  3. (29)
+    In which country do you currently reside? [Open Text - Likely]
+  4. (30)
+    What are the main languages you speak daily? Select up to three languages. [Open Text]
+  5. (31)
+    Did you experience any barriers accessing ChatGPT or submitting feedback about ChatGPT?
+    • Computer hardware barriers
+    • Internet connectivity barriers
+    • Health and medical barriers
+    • Language barriers
+    • Other (Specify)
+    • None of the above
+  6. (32)
+    How often do you engage with or read about AI-related topics? e.g., news articles, research papers, podcasts, and so on.
+    • Daily
+    • Weekly
+    • Monthly
+    • Rarely
+    • Never
+  7. (33)
+    Before using ChatGPT, have you used any AI, machine learning, or data science tools, beyond basic interaction or exploration?
+Please select “Yes” if you have used AI, machine learning, or data science tools for a project, work, or in-depth exploration.
+    • Yes
+    • No
+  8. (34)
+    Have you completed any formal education, training, or coursework related to AI, machine learning, or data science?
+    • Yes
+    • No
+  9. (35)
+    Are you currently working in, or have you ever worked in, a field related to AI, machine learning, or data science?
+    • Yes
+    • No
+  10. (36)
+    How much do you agree with the following statement: “I feel confident in my ability to evaluate and provide feedback on the performance and/or impact of generative A systems, such as ChatGPT”?
+    • Strongly agree
+    • Somewhat agree
+    • Neither agree nor disagree
+    • Somewhat disagree
+    • Strongly disagree
+Appendix B Survey Results
+Appendix B compiles the survey results from this study.
 Table 9. Frequency of ChatGPT Usage
 Frequency Count Percentage
+Daily 64 12.2%
 A few times a week 216 41.1%
@@
 Rarely 65 12.4%
-Daily 64 12.2%
 Total 525 100.0%
-(Response rate: 525/526)
+(Response rate: 525/526).
 Table 10. Ways of Using ChatGPT
@@
 Other 22 4.2%
-(Response rate: 526/526, Multiple selections allowed)
+(Response rate: 526/526, Multiple selections allowed).
 Table 11. Submitted Thumbs-Up/Down Feedback
@@
 Total 526 100.0%
-(Response rate: 526/526)
+(Response rate: 526/526).
 Table 12. Frequency of Thumbs-Up/Down Feedback
 Frequency Count Percentage
+Every time I have used ChatGPT 23 5.8%
+Most times I have used ChatGPT 92 23.0%
+About half the time I have used ChatGPT 98 24.5%
 Sometimes I have used ChatGPT 139 34.8%
-About half the time I have used ChatGPT 98 24.5%
-Most times I have used ChatGPT 92 23.0%
 Rarely 48 12.0%
-Every time I have used ChatGPT 23 5.8%
 Total 400 100.0%
-(Response rate: 400/400)
+(Response rate: 400/400).
 Table 13. Submitted Open-Text Feedback
 Submitted feedback Count Percentage
+Yes 152 29.0%
 No 248 47.1%
-Yes 152 29.0%
 N/A 126 23.9%
@@
 Submitted feedback Count Percentage
+Yes 152 38.0%
 No 248 62.0%
-Yes 152 38.0%
 Total 400 100.0%
-(Response rate: 400/400)
+(Response rate: 400/400).
 Table 15. Frequency of Open-Text Feedback
 Frequency Count Percentage
+Every time I have used ChatGPT 3 2.0%
+Most times I have used ChatGPT 20 13.2%
+About half the time I have used ChatGPT 17 11.2%
 Sometimes I have used ChatGPT 77 50.7%
 Rarely 35 23.0%
-Most times I have used ChatGPT 20 13.2%
-About half the time I have used ChatGPT 17 11.2%
-Every time I have used ChatGPT 3 2.0%
 Total 152 100.0%
-(Response rate: 152/152)
+(Response rate: 152/152).
 Table 16. Submitted Feedback After Regenerating a Response
@@
 Total 526 100.0%
-(Response rate: 526/526)
+(Response rate: 526/526).
 Table 17. Frequency of Feedback After Regenerating a Response
 Frequency Count Percentage
-Sometimes I have regenerated a response 120 36.3%
+Every time I have regenerated a response 25 7.6%
 Most times I have regenerated a response 67 20.2%
 About half the time I have regenerated a response 65 19.6%
+Sometimes I have regenerated a response 120 36.3%
 Rarely 54 16.3%
-Every time I have regenerated a response 25 7.6%
 Total 331 100.0%
-(Response rate: 331/331)
+(Response rate: 331/331).
 Table 18. Reasons for Not Submitting Feedback
@@
 Not enough time or interest in submitting feedback 21 26.9%
-Satisfied with the performance and have not felt the need to submit feedback 17 21.8%
+Satisfied with the performance and have not felt need to submit feedback 17 21.8%
 Unsure how to provide feedback 7 9.0%
@@
 Total 78 100.0%
-(Response rate: 78/78)
+(Response rate: 78/78).
 Table 19. Motivations for Providing Feedback
@@
 Other 15
-(Response rate: 152/152)
+(Response rate: 152/152).
 Note: Weighted frequencies based on a two-part question. Part A (multiple selection) responses weighted as 1; Part B (main motivation) responses weighted as 2.
@@
 Response Count Percentage
+Yes 17 11.2%
 No 135 88.8%
-Yes 17 11.2%
 Total 152 100.0%
-(Response rate: 152/152)
+(Response rate: 152/152).
 Table 21. Satisfaction with Follow-up
 Satisfaction level Count Percentage
+Extremely satisfied 5 29.4%
 Somewhat satisfied 10 58.8%
-Extremely satisfied 5 29.4%
 Neither satisfied nor dissatisfied 2 11.8%
+Somewhat dissatisfied 0 0%
+Extremely dissatisfied 0 0%
 Total 17 100.0%
-(Response rate: 17/17)
+(Response rate: 17/17).
 Table 22. Reasons for Submitting Feedback Addressed
 Agreement level Count Percentage
+Strongly agree 8 5.3%
+Somewhat agree 36 23.7%
 Neither agree nor disagree 72 47.4%
-Somewhat agree 36 23.7%
 Somewhat disagree 25 16.4%
 Strongly disagree 11 7.2%
-Strongly agree 8 5.3%
 Total 152 100.0%
-(Response rate: 152/152)
+(Response rate: 152/152).
 Table 23. Ease of Submitting Feedback
@@
 Somewhat difficult 4 0.9%
+Extremely difficult 0 0%
 Total 442 100.0%
-(Response rate: 442/443)
+(Response rate: 442/443).
 Table 24. Likelihood of Submitting Feedback in Future
 Likelihood Count Percentage
+Extremely likely 161 30.7%
 Somewhat likely 255 48.7%
-Extremely likely 161 30.7%
 Neither likely nor unlikely 67 12.8%
@@
 Total 524 100.0%
-(Response rate: 524/526)
+(Response rate: 524/526).
 Table 25. Provided Feedback via Other Platforms or Channels
 Provided feedback Count Percentage
+Yes 103 19.6%
 No 423 80.4%
-Yes 103 19.6%
 Total 526 100.0%
-(Response rate: 526/526)
+(Response rate: 526/526).
 Table 26. Other Channels or Platforms
@@
 LinkedIn 10 9.7%
-Contacting OpenAI (e.g., via email) 10 9.7%
+Contacting OpenAI (e.g., via e-mail) 10 9.7%
 Other online discussion forums 9 8.7%
 Mastodon 3 2.9%
-(Response rate 103/103, Multiple selections allowed)
+(Response rate 103/103, Multiple selections allowed).
 Table 27. Likelihood of Submitting Feedback via Other Channels or Platforms in Future
 Likelihood Count Percentage
+Extremely likely 28 27.2%
 Somewhat likely 46 44.7%
-Extremely likely 28 27.2%
 Neither likely nor unlikely 15 14.6%
@@
 Total 103 100.0%
-(Response rate: 103/103)
+(Response rate: 103/103).
 Table 28. Age
@@
 Total 524 100.0%
-(Response rate: 524/526)
+(Response rate: 524/526).
 Table 29. Gender
@@
 Total 521 100.0%
-(Response rate: 521/526)
+(Response rate: 521/526).
 Table 30. Region
@@
 Total 507 100.0%
-(Response rate: 507/526)
+(Response rate: 507/526).
 Table 31. Languages Spoken Daily (Grouped)
@@
 Total 509 100.0%
-(Response rate: 509/526, up to 3 selections allowed)
-Notes: ‘Other African languages’ includes Setswana, Xhosa, Yoruba, Swahili, and Hausa; ’Other European languages’ includes Finnish, Hungarian, Latvian, Russian, Norwegian, Danish, Swedish, Catalan, Ukrainian, and Lithuanian; ’Other Asian languages’ includes Hindi, Hebrew, Turkish, Arabic, Mongolian, Korean, Japanese, Persian, Bengali, Tagalog, Thai, and Sinhalese; ’Other languages’ includes Portuguese (Brazilian), Slovenian, Croatian, and unspecified others.
+(Response rate: 509/526, up to 3 selections allowed).
+Notes: “Other African languages” includes Setswana, Xhosa, Yoruba, Swahili, and Hausa; “Other European languages” includes Finnish, Hungarian, Latvian, Russian, Norwegian, Danish, Swedish, Catalan, Ukrainian, and Lithuanian; “Other Asian languages” includes Hindi, Hebrew, Turkish, Arabic, Mongolian, Korean, Japanese, Persian, Bengali, Tagalog, Thai, and Sinhalese; “Other languages” includes Portuguese (Brazilian), Slovenian, Croatian, and unspecified others.
 Table 32. Barriers Accessing ChatGPT or Submitting Feedback
 Barriers Count Percentage
+Barriers 118 22.7%
 No barriers 401 77.3%
-Barriers 118 22.7%
 Total 519 100.0%
-(Response rate: 519/526)
+(Response rate: 519/526).
 Table 33. Frequency of Engaging with AI-related Topics
 Frequency Count Percentage
+Daily 87 16.6%
 Weekly 243 46.5%
 Monthly 99 18.9%
-Daily 87 16.6%
 Rarely 85 16.3%
@@
 Total 523 100.0%
-(Response rate: 523/526)
+(Response rate: 523/526).
 Table 34. Prior Use of AI, ML, or Data Science Tools
 Response Count Percentage
+Yes 214 41.0%
 No 308 59.0%
-Yes 214 41.0%
 Total 522 100.0%
-(Response rate: 522/526)
+(Response rate: 522/526).
 Table 35. Completed Education or Training in AI, ML, or Data Science
 Response Count Percentage
+Yes 138 26.5%
 No 382 73.5%
-Yes 138 26.5%
 Total 520 100.0%
-(Response rate: 520/526)
+(Response rate: 520/526).
 Table 36. Currently or Previously Working in AI, ML, or Data Science
 Response Count Percentage
+Yes 111 21.3%
 No 410 78.7%
-Yes 111 21.3%
 Total 521 100.0%
-(Response rate: 521/526)
+(Response rate: 521/526).
 Table 37. Confidence Submitting Feedback
 Confidence level Count Percentage
+Strongly agree 114 21.8%
 Somewhat agree 296 56.5%
-Strongly agree 114 21.8%
 Neither agree nor disagree 79 15.1%
@@
 Total 524 100.0%
-(Response rate: 524/526)
-Appendix B Statistical Relationships and Test Results
-Appendix B presents our statistical analysis of relationships between key variables in our study.
-Table 31 presents the results of chi-square tests of associations between various factors such as user characteristics, feedback behaviours, and perceptions of submitting feedback through the ChatGPT interface.
-Table 31. Chi-square Test Results
+(Response rate: 524/526).
+Appendix C Statistical Relationships and Test Results
+Appendix C presents the statistical analysis of relationships between key variables for this study.
+Table 38. Chi-squared Test Results
 Relationship $\chi^{2}$ $p$ Adj. $p$ df CV Interp. $n$
-Thumbs-Up/down & Regenerating feedback 50.419 ¡.001 ¡.001* 1 .319 Mod. 497
-Education/Training in AI & Confidence 12.773 .002 .023* 2 .160 Weak 497
-Working in AI & Barriers 9.821 .002 .023* 1 .143 Weak 477
-Region & Open-text feedback 13.990 .003 .023* 3 .192 Mod. 378
-Prior AI use & Open-text feedback 8.926 .003 .023* 1 .154 Weak 378
-Thumbs-Up/down & ChatGPT frequency 13.689 .008 .045* 4 .166 Weak 497
-AI topics frequency & Confidence 20.726 .008 .045* 8 .144 Weak 497
-Age & Confidence 17.068 .009 .045* 6 .131 Weak 497
-Education/Training in AI & Open-text 6.132 .013 .059 1 .127 Weak 378
-Gender & Regenerating feedback 10.427 .015 .061 3 .145 Weak 497
-Age & Barriers 9.372 .025 .078 3 .140 Weak 477
-Open-text & ChatGPT frequency 10.956 .027 .078 4 .170 Mod. 378
-AI topics frequency & Open-text 10.937 .027 .078 4 .170 Mod. 378
-Region & Regenerating feedback 9.151 .027 .078 3 .136 Weak 497
-Working in AI & Open-text feedback 4.660 .031 .080 1 .111 Weak 378
-Age & Regenerating feedback 8.678 .034 .080 3 .132 Weak 497
-Prior AI use & Barriers 4.512 .034 .080 1 .097 Negl. 477
-Education/Training in AI & Barriers 4.010 .045 .101 1 .092 Negl. 477
-Gender & Confidence 12.513 .051 .103 6 .112 Weak 497
-Prior AI use & Confidence 5.953 .051 .103 2 .109 Weak 497
-Working in AI & Confidence 5.632 .060 .114 2 .106 Weak 497
-Open-text & Regenerating feedback 3.424 .064 .117 1 .095 Negl. 378
-AI topics frequency & Thumbs-Up/down 7.615 .107 .186 4 .124 Weak 497
-Working in AI & Regenerating feedback 2.375 .123 .205 1 .069 Negl. 497
-Regenerating & ChatGPT frequency 5.004 .287 .459 4 .100 Weak 497
-Age & Open-text feedback 3.342 .342 .526 3 .094 Weak 378
-Gender & Thumbs-Up/down 3.160 .368 .545 3 .080 Weak 497
-AI topics frequency & Barriers 3.962 .411 .554 4 .091 Weak 477
-AI topics frequency & Regenerating 3.932 .415 .554 4 .089 Weak 497
-Prior AI use & Thumbs-Up/down 0.666 .415 .554 1 .037 Negl. 497
-Gender & Barriers 2.482 .478 .607 3 .072 Weak 477
-Education/Training in AI & Thumbs 0.468 .494 .607 1 .031 Negl. 497
-Working in AI & Thumbs-Up/down 0.453 .501 .607 1 .030 Negl. 497
-Gender & Open-text feedback 1.801 .615 .723 3 .069 Weak 378
-Age & Thumbs-Up/down 1.713 .634 .725 3 .059 Negl. 497
-Prior AI use & Regenerating feedback 0.140 .708 .773 1 .017 Negl. 497
-Region & Barriers 1.361 .715 .773 3 .053 Negl. 477
-Education/Training in AI & Regenerating 0.065 .799 .841 1 .011 Negl. 497
-Region & Thumbs-Up/down 0.708 .871 .894 3 .038 Negl. 497
-Region & Confidence 1.548 .956 .956 6 .039 Negl. 497
+Thumbs-Up/down and After Regenerating 50.419 ¡.001 ¡.001* 1 .319 Mod. 497
+Education/Training in AI and Confidence 12.773 .002 .023* 2 .160 Weak 497
+Working in AI and Barriers 9.821 .002 .023* 1 .143 Weak 477
+Region and Open-text feedback 13.990 .003 .023* 3 .192 Mod. 378
+Prior AI use and Open-text feedback 8.926 .003 .023* 1 .154 Weak 378
+Thumbs-Up/down and ChatGPT frequency 13.689 .008 .045* 4 .166 Weak 497
+AI topics frequency and Confidence 20.726 .008 .045* 8 .144 Weak 497
+Age and Confidence 17.068 .009 .045* 6 .131 Weak 497
+Education/Training in AI and Open-text 6.132 .013 .059 1 .127 Weak 378
+Gender and After Regenerating 10.427 .015 .061 3 .145 Weak 497
+Age and Barriers 9.372 .025 .078 3 .140 Weak 477
+Open-text and ChatGPT frequency 10.956 .027 .078 4 .170 Mod. 378
+AI topics frequency and Open-text 10.937 .027 .078 4 .170 Mod. 378
+Region and After Regenerating 9.151 .027 .078 3 .136 Weak 497
+Working in AI and Open-text feedback 4.660 .031 .080 1 .111 Weak 378
+Age and After Regenerating 8.678 .034 .080 3 .132 Weak 497
+Prior AI use and Barriers 4.512 .034 .080 1 .097 Negl. 477
+Education/Training in AI and Barriers 4.010 .045 .101 1 .092 Negl. 477
+Gender and Confidence 12.513 .051 .103 6 .112 Weak 497
+Prior AI use and Confidence 5.953 .051 .103 2 .109 Weak 497
+Working in AI and Confidence 5.632 .060 .114 2 .106 Weak 497
+Open-text and After Regenerating 3.424 .064 .117 1 .095 Negl. 378
+AI topics frequency and Thumbs-Up/down 7.615 .107 .186 4 .124 Weak 497
+Working in AI and After Regenerating 2.375 .123 .205 1 .069 Negl. 497
+Regenerating and ChatGPT frequency 5.004 .287 .459 4 .100 Weak 497
+Age and Open-text feedback 3.342 .342 .526 3 .094 Weak 378
+Gender and Thumbs-Up/down 3.160 .368 .545 3 .080 Weak 497
+AI topics frequency and Barriers 3.962 .411 .554 4 .091 Weak 477
+AI topics frequency and Regenerating 3.932 .415 .554 4 .089 Weak 497
+Prior AI use and Thumbs-Up/down 0.666 .415 .554 1 .037 Negl. 497
+Gender and Barriers 2.482 .478 .607 3 .072 Weak 477
+Education/Training in AI and Thumbs 0.468 .494 .607 1 .031 Negl. 497
+Working in AI and Thumbs-Up/down 0.453 .501 .607 1 .030 Negl. 497
+Table 38. Continued
+Relationship $\chi^{2}$ $p$ Adj. $p$ df CV Interp. $n$
+Gender and Open-text feedback 1.801 .615 .723 3 .069 Weak 378
+Age and Thumbs-Up/down 1.713 .634 .725 3 .059 Negl. 497
+Prior AI use and After Regenerating 0.140 .708 .773 1 .017 Negl. 497
+Region and Barriers 1.361 .715 .773 3 .053 Negl. 477
+Education/Training in AI and Regenerating 0.065 .799 .841 1 .011 Negl. 497
+Region and Thumbs-Up/down 0.708 .871 .894 3 .038 Negl. 497
+Region and Confidence 1.548 .956 .956 6 .039 Negl. 497
 Notes:
@@
   • * Indicates statistically significant $p$-value after adjustment.
-  • $\chi^{2}$ test was not performed between Thumbs-Up/down & Open-text feedback.
+  • $\chi^{2}$ test was not performed between Thumbs-Up/down and Open-text feedback.
 Open-text feedback was only available to users who provided Thumbs-Up/down feedback, making these variables dependent.
-  • Table 32 provides a guide for interpreting Cramér’s V values, in accordance with (Cohen, 2013).
+  • Table 39 provides a guide for interpreting Cramér’s V values, in accordance with (Cohen, 2013). Thresholds for df $\geq$ 3 are approximate guidelines.
 Table 39. Interpretation of Cramér’s V